### v1

# Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization

Abstract
Response-adaptive randomization (RAR) methods use accumulated data to adapt randomization probabilities, aiming to increase the probability of allocating patients to effective treatments. A popular RAR method is Thompson sampling, which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective. However, its high variability early in a trial can also increase the risk of assigning patients to inferior treatments. We propose a principled method based on Bayesian hypothesis testing to mitigate this issue. Specifically, we introduce a point null hypothesis that postulates equal effectiveness of treatments. This induces shrinkage toward equal randomization probabilities, with the degree of shrinkage controlled by the prior probability of the null hypothesis. Equal randomization and Thompson sampling arise as special cases when the prior probability is set to one or zero, respectively. Simulated and real-world examples illustrate that the proposed method balances highly variable Thompson sampling with static equal randomization. A simulation study demonstrates that the method can mitigate issues with ordinary Thompson sampling and has comparable statistical properties to Thompson sampling with common ad hoc modifications such as power transformation and probability capping. We implement the method in the open-source R package brar, enabling experimenters to easily perform point null Bayesian RAR and support more effective randomization of patients.
Keywords: Adaptive trials, A/B testing, Bayes factor, Bayesian model averaging, sharp null hypothesis, spike-and-slab prior
1 Introduction
Response-adaptive randomization (RAR) methods, sometimes also called outcome-adaptive randomization methods, randomly allocate experimental units (e.g., patients or animals) to treatments in a manner that is informed by accumulating data (Thall and Wathen,, 2007; Berry et al.,, 2010; Grieve,, 2016; Robertson et al.,, 2023). A popular approach is Thompson sampling (Thompson,, 1933), which randomizes participants proportionally to the Bayesian posterior probability that each treatment is the most effective. Such RAR methods are attractive because they naturally balance gathering information on treatment effectiveness and assigning subjects to effective treatments.
Despite its benefits, there are also various challenges with Thompson sampling. For example, the method can exhibit high variability, particularly in the early stages of a study when posterior uncertainty is high (Thall et al.,, 2015). This can lead to erratic allocation probabilities, ethical concerns (e.g., exposing participants to inferior treatments with high probability), and inferential challenges (e.g., reduction of statistical power or biased effect estimates). Consequently, there has been substantial interest in modifying Thompson sampling to make it more reliable.
For example, Thall and Wathen, (2007) propose to use the randomization probability $\pi=p^{c}/\{p^{c}+(1-p)^{c}\}$ where $p$ is the posterior probability that the experimental treatment is more effective than the control treatment, and $c$ is an additional parameter that controls the variability of the method. Setting $c=1$ produces Thompson sampling, while $c<1$ reduces variability with $c=1/2$ being often recommended. Another approach to reduce extreme randomization probabilities is to cap them, for instance, setting them to 10% or 90% if a method assigns more extreme probabilities (Thall and Wathen,, 2007; Lee and Lee,, 2021). Finally, RAR methods are often combined with “burn-in” periods at the start of the study, during which units are randomized with equal probabilities to mitigate high variability (Thall and Wathen,, 2007; Wathen and Thall,, 2017; Robertson et al.,, 2023).
While such ad hoc modifications can address some of the limitations of RAR methods, they conflict with the principles of coherent Bayesian learning. For example, a transformed or capped posterior probability does no longer correspond to an actual posterior probability, and it cannot be used as a genuine prior for future data. This raises the question of whether it is possible to devise a RAR method with desirable properties, such as reduced variability compared to Thompson sampling, that is coherent with Bayesian principles, and if so, how it relates to these ad hoc modifications.
In this paper, we propose a novel RAR method that reduces variability in a coherent Bayesian manner, which we term “point null Bayesian RAR”. The idea is to consider a point null hypothesis postulating that treatments are equally effective. This is equivalent to using a “spike-and-slab” prior (Raftery et al.,, 1997), also known as “lump-and-smear” prior, which is a mixture of a point mass at equal effectiveness and a probability density elsewhere (Spiegelhalter et al.,, 2004, Section 5.5.4). The prior probability of the point null hypothesis determines the mixture weight. As we will show, setting this prior probability to zero produces equal randomization, whereas setting it to one produces Thompson sampling. The proposed method thus interpolates between equal randomization and Thompson sampling in a coherent Bayesian way. As a by-product, posterior probabilities and Bayes factors are also obtained. These can be used to monitor evidence of the effectiveness of each treatment in a manner that is aligned with the randomization probabilities.
In the following Section 2 we introduce the general idea of the method in more detail, followed by tailoring it to the setting of approximately normal effect estimates (Section 3) and binary outcomes (Section 4). In Section 5, we then illustrate the method on data from the ECMO trial (Bartlett et al.,, 1985), followed by evaluating its statistical properties in a simulation study (Section 6). The paper ends with concluding remarks on advantages, limitations, and opportunities for future research (Section 7). Appendix A illustrates our R package brar for performing point null Bayesian RAR. Appendix B provides further details on our simulation study.
2 Point null Bayesian RAR
We now explain the general idea of point null Bayesian RAR, without going into specifics such as data distribution or computation (these will follow in the subsequent sections). Throughout we will assume that we have observed data $y$ and we want to use these to randomize a future experimental unit. We start with the basic but important setting of one control and one treatment group, and extend it afterwards to multiple treatment groups.
2.1 Two group comparisons
In case there is a control group and only one treatment group, we consider the hypotheses:
$H_{-}\colon\text{Treatment is less effective than control}$
$H_{0}\colon\text{Treatment and control are equally effective}$
$H_{+}\colon\text{Treatment is more effective than control}$
How exactly these statements are translated into statistical hypotheses related to parameters depends on the type of data and model used, but often relates to an effect size parameter being less, equal, or greater than zero. There may also be situations where there is no control group but only two competing treatments. In this case, the method detailed here is still applicable but with the control group replaced by a reference group (the choice of the reference may be somewhat arbitrary).
The Bayesian posterior probability of a hypothesis $H_{i}\in\{H_{-},H_{0},H_{+}\}$ can then be calculated by
$\Pr(H_{i}\mid y)=\frac{p(y\mid H_{i})\Pr(H_{i})}{\sum\limits_{j\in\{-,0,+\}}p(y\mid H_{j})\Pr(H_{j})}=\left\{\sum_{j\in\{-,0,+\}}\mathrm{BF}_{ji}(y)\,\frac{\Pr(H_{j})}{\Pr(H_{i})}\right\}^{-1}$ (1)
where $\Pr(H_{i})$ is the prior probability of hypothesis $H_{i}$,
$p(y\mid H_{i})=\int_{\Theta}p(y\mid\theta)\,p(\theta\mid H_{i})\,\mathrm{d}\theta$
is the marginal likelihood of the data $y$ under $H_{i}$ obtained from marginalizing the likelihood $p(y\mid\theta)$ with respect to the prior distribution $p(\theta\mid H_{i})$ assigned to the model parameters $\theta\in\Theta$ under $H_{i}$, and
$\mathrm{BF}_{ji}(y)=\frac{\Pr(H_{j}\mid y)}{\Pr(H_{i}\mid y)}\,\bigg/\,\frac{\Pr(H_{j})}{\Pr(H_{i})}=\frac{p(y\mid H_{j})}{p(y\mid H_{i})}$ (2)
is the Bayes factor contrasting $H_{j}$ to $H_{i}$ (Jeffreys,, 1939; Good,, 1958). The Bayes factor (2) is the updating factor of the prior odds of $H_{j}$ to $H_{i}$ to the corresponding posterior odds (first equality), which is equivalent to the ratio of marginal likelihoods of the data under $H_{j}$ and $H_{i}$ (second equality). The posterior probabilities (1) can thus be computed from the marginal likelihoods of the data under each considered hypotheses along with their prior probabilities, or from the set of Bayes factors and prior hypothesis odds relative to some reference hypothesis (Kass and Raftery,, 1995).
Regardless in which way they are computed, the question is how to translate posterior probabilities into randomization probabilities. We propose to randomize a future unit to the treatment group with probability
$\pi=\Pr(H_{+}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2},$ (3)
while the probability to randomize to the control group is consequently
$1-\pi=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2}.$ (4)
Figure 1 shows such randomization probabilities for different combinations of posterior probabilities. We can see that the randomization probability $\pi$ shrinks towards 50% as the posterior probability of the null hypothesis $\Pr(H_{0}\mid y)$ increases. For example, if $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$, as indicated by the black asterisk in Figure 1, the probability to randomize to the treatment group is $\pi=0.2+0.5/2=45\%$.
Figure 1: Ternary plot where the color of each point depicts the treatment randomization probability $\pi$ for a particular combination of posterior probabilities of $H_{+}$ (bottom axis), $H_{-}$ (left axis), and $H_{0}$ (right axis). For example, the black asterisk denotes the combination where $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$ with corresponding randomization probability $\pi=45\%$.
This scheme can be motivated as Bayesian hypothesis averaged randomization probability where the hypothesis-specific treatment randomization probabilities are $0\%$, $50\%$, and $100\%$, under $H_{-}$, $H_{0}$, and $H_{+}$, respectively. That is, if we would know that $H_{+}$ or $H_{-}$ is true, we should assign the next patient to the treatment or control group, respectively, with probability one to maximize utility (patient benefit). On the other hand, if we knew that $H_{0}$ is true, then randomizing the assignment with $\pi=50\%$ seems sensible.
Interestingly, the randomization scheme (3) reduces to Thompson sampling when the prior probability of $H_{0}$ is set to zero since then $\Pr(H_{0}\mid y)=0$ regardless of the data, and consequently $\pi=\Pr(H_{+}\mid y)$. However, the scheme induces shrinkage toward randomization probabilities of $\pi=50\%$ otherwise. In the most extreme case when $\Pr(H_{0})=1$, equal randomization is obtained as then $\Pr(H_{0}\mid y)=1$ regardless of the data, and consequently $\pi=50\%$. The scheme thus interpolates between Thompson sampling and equal randomization in a coherent Bayesian way.
2.2 More than two groups
Suppose there are $K>1$ treatment groups in addition to the control group. In this case, we may modify the procedure and consider the hypotheses:
$H_{-}\colon\text{All treatments are less effective than control}$
$H_{0}\colon\text{All treatments are equally effective as control}$
[…]
$\vdots$
[…]
Posterior probabilities of each hypothesis can be computed from the marginal likelihoods and prior hypotheses probabilities with the summation in (1) extended to encompass all hypotheses (i.e., summing over $j\in\{-,0,+1,\dots,+K\}$). Similarly, they can be translated into randomization probabilities
$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1}$ (5)
with corresponding control randomization probability
$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1},$ (6)
which reduce to the randomization probabilities (3) and (4) for $K=1$.
Also in the multi-treatment case, the randomization probabilities are shrunken towards equal randomization $\pi=1/(K+1)$ by introducing the null hypothesis $H_{0}$. Similarly, Thompson sampling and equal randomization are obtained by setting $\Pr(H_{0})=0$ and $\Pr(H_{0})=1$, respectively, since then the posteriors $\Pr(H_{0}\mid y)=0$ and $\Pr(H_{0}\mid y)=1$ are obtained for any observed data $y$.
In this paper, we focus on schemes (5) and (6), which induce shrinkage toward equal randomization. However, it is important to note that there are alternatives to this approach. For example, it may be desired to shrink to different “baseline” randomization probabilities than equal randomization probabilities. This can be achieved by modifying the multiplicative factor of $\Pr(H_{0}\mid y)$ in (5) from $1/(K+1)$ to the desired baseline randomization probability. For example, if the goal is to minimize the standard errors of the treatment effect estimates by using $\sqrt{K}:1:\dots:1$ square-root allocation of the control to treatments (Dunnett,, 1955), we may use
$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+\sqrt{K}}$
which leads to the control randomization probability
$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{\sqrt{K}}{K+\sqrt{K}}$
These correctly shrink RAR probabilitites towards the Dunnett-type randomization probabilities, $\pi_{i}=1/(K+\sqrt{K})$ for treatment $i=1,\dots,K$, and $1-\sum_{i=1}^{K}\pi_{i}=\sqrt{K}/(K+\sqrt{K})$ for control.
3 Point null Bayesian RAR under approximate normality
We will expand on point null Bayesian RAR in the setting where the data are summarized by an asymptotically normally distributed effect estimate. This does not mean that the raw data (e.g., a vector of outcomes and a matrix of covariates), from which the estimate is computed, need to be normally distributed. For instance, the regression coefficients from generalized linear models estimated using maximum likelihood satisfy asymptotic normality even if the data are themselves not normally distributed. This is useful because it allows us to efficiently compute posterior and randomization probabilities without the need for simulation. Although this framework is widely applicable, improvements may be possible by considering the exact distribution of the data. This will be detailed for binary outcomes in Section 4.
3.1 Two group comparisons
Assume that the data are summarized by $y=\{\hat{\theta},\sigma\}$, where $\hat{\theta}$ is an effect estimate of the true treatment effect $\theta$ that quantifies the effect of a treatment over the control and $\sigma$ is the standard error of the estimate. For example, $\hat{\theta}$ could be an estimated (standardized) mean difference, log odds/rate/hazard ratio, risk difference, or regression coefficient. Suppose further that the estimate is (at least approximately) normally distributed, i.e., $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2})$.
If the effect $\theta$ is oriented such that a positive effect indicates treatment benefit, the three hypotheses from Section 2.1 translate into:
$H_{-}\colon\theta<0$ versus $H_{0}\colon\theta=0$ versus $H_{+}\colon\theta>0$
The point null hypothesis $H_{0}$ is a simple hypothesis with no free parameters, or equivalently, a point (Dirac) prior at zero. The $H_{-}$ and $H_{+}$ hypotheses are composite hypotheses and require prior distributions for the effect $\theta$. A natural choice is a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ normal distribution whose support is truncated to the negative and positive side, respectively. This distribution may be specified based on prior knowledge (e.g., previous studies). In the absence of prior knowledge, it seems sensible to center the prior on zero ($\mu=0$) to represent clinical equipoise (Freedman,, 1987), as a zero-centered prior gives equal probability to harmful and beneficial effects. Averaging the prior over the three hypotheses leads to a spike-and-slab prior, as illustrated in Figure 2.
Figure 2: Illustration of spike-and-slab prior for the effect $\theta$. A point prior at 0 is assumed under $H_{0}$. A normal prior $\theta\sim\mathrm{N}(0,1)$ with support truncated to the positive or negative side is assumed under $H_{+}$ and $H_{-}$, respectively. These priors are averaged assuming prior hypothesis probabilities $\Pr(H_{0})=0.5$, $\Pr(H_{+})=0.25$, and $\Pr(H_{-})=0.25$.
To compute posterior hypothesis probabilities, specification of prior hypothesis probabilities is required. The prior probability of the null hypothesis $\Pr(H_{0})$ represents the a priori plausibility of an absent effect, and also acts as a tuning parameter that controls the degree of variability of RAR. An intuitive default is $\Pr(H_{0})=0.5$ as it represents the equipoise position of equal probability of an absent effect relative to a present (either harmful or beneficial) effect (Johnson,, 2013). For a given $\Pr(H_{0})$, it is then natural to set the prior probabilities of the other two hypotheses to $\Pr(H_{+})=\{1-\Pr(H_{0})\}\times\Phi(\mu/\tau)$ and $\Pr(H_{-})=\{1-\Pr(H_{0})\}\times\Phi(-\mu/\tau)$ so that to the prior distribution averaged over $H_{-}$ and $H_{+}$ is again the $\mathrm{N}(\mu,\tau^{2})$ normal distribution that was truncated in the first place. For example, when setting $\Pr(H_{0})=0.5$ and specifying a zero-centered prior ($\mu=0$) as in Figure 2, we obtain $\Pr(H_{+})=\Pr(H_{-})=0.5\times 0.5=0.25.$
With the prior densities and prior hypothesis probabilities specified, we can compute the marginal likelihood of the observed effect estimate under each hypothesis, and in turn obtain posterior probabilities (1). In the conjugate normal likelihood and prior framework, all of them can be straightforwardly derived in closed-form. Denoting by $\mathrm{N}(x\mid m,v)$ the normal density function with mean $m$ and variance $v$ evaluated at $x$, the marginal likelihoods are given by
$p(\hat{\theta}\mid H_{-})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(-\mu/\tau)}$ (7a)
$p(\hat{\theta}\mid H_{0})$ $=\mathrm{N}(\hat{\theta}\mid 0,\sigma^{2})$ (7b)
$p(\hat{\theta}\mid H_{+})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(\mu_{*}/\tau_{*})}{\Phi(\mu/\tau)}$ (7c)
with posterior mean and variance
$\mu_{*}=\frac{\hat{\theta}/\sigma^{2}+\mu/\tau^{2}}{1/\sigma^{2}+1/\tau^{2}}$ and $\tau_{*}^{2}=\frac{1}{1/\sigma^{2}+1/\tau^{2}}.$
Taking ratios of marginal likelihoods produces the Bayes factors
[…]
and the Bayes factors for other hypothesis comparisons can be obtained by transitivity and reciprocity, for example, $\mathrm{BF}_{-0}=\mathrm{BF}_{+0}\,/\,\mathrm{BF}_{+-}$. Posterior probabilities can now be obtained by plugging the Bayes factors and prior odds into (1), leading to
[…]
These Bayes factors and posterior probabilities can be monitored as data accumulate to see how the evidence for the hypotheses change. Moreover, they can be used to define the randomization probabilities via (3) leading to
$\pi$ $=\left(1+\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(\mu_{*}/\tau_{*})}+\frac{\Pr(H_{0})}{1-\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{\hat{\theta}^{2}}{\sigma^{2}}-\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}\right\}\right]\,\times\,\frac{\sqrt{1+\tau^{2}/\sigma^{2}}}{\Phi(\mu_{*}/\tau_{*})}\right)^{-1}$
$\phantom{=}+\left(1+\frac{1-\Pr(H_{0})}{\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}-\frac{\hat{\theta}^{2}}{\sigma^{2}}\right\}\right]\,\bigg/\,\sqrt{1+\frac{\tau^{2}}{\sigma^{2}}}\right)^{-1}\,\bigg/\,2.$ (8)
As expected, the randomization probability (3.1) approaches equal randomization as the prior probability of $H_{0}$ increases to one (i.e., $\pi\to 50\%$ as $\Pr(H_{0})\nearrow 1$), while it approaches the ordinary Bayesian posterior tail probability of $\theta>0$ based on a normal prior $\theta\sim\mathrm{N}(\mu,\tau^{2})$ as the prior probability of $H_{0}$ decreases to zero (i.e., $\pi\to\Phi(\mu_{*}/\tau_{*})$ as $\Pr(H_{0})\searrow 0$).
Figure 3: Evolution of Bayesian RAR probabilities for one treatment and control group. In each step one observation from the control and one from the treatment group are simulated from a normal distribution with a true standard deviation of 1 and assuming a true mean difference $\theta=0.25$. Randomization probabilities are then computed assuming a normal spike-and-slab prior centered at zero with standard deviation $\tau=1$, and different prior probabilities $\Pr(H_{0})$.
Figure 3 shows sequences of randomization probabilities computed from simulated normal data. To enable comparison of randomization probabilities across different prior probabilities $\Pr(H_{0})$, the data were not simulated under RAR but by simulating an additional observation from the treatment and control groups at each step. We can see that the probability to randomize to the treatment group based on $\Pr(H_{0})=1$ remains static at $50\%$ (yellow line). In contrast, the randomization probabilities based on $\Pr(H_{0})<1$ tend towards $100\%$ as more data accumulate, as expected, given that the data were simulated under a beneficial treatment effect. Moreover, a clear ordering is visible: Probabilities under $\Pr(H_{0})=0$, which corresponds to Thompson sampling, are the most extreme and may even go strongly in the “wrong” direction. For example, the probability to randomize to control is higher than 75% at some point during the earlier stages of the study. In contrast, probabilities based on $0<\Pr(H_{0})<1$ show the same qualitative behavior but are less extreme. Setting a higher prior probability $\Pr(H_{0})$ thus reduces the variability of randomization probabilities but also makes convergence to a probability of 100% for the more effective treatment slower.
3.2 More than two groups
Suppose now there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to the control. A natural generalization is to stack the estimates into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume an approximate $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\mid\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the covariance matrix of $\boldsymbol{\hat{\theta}}$. For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ the corresponding covariance matrix.
The hypotheses from Section 2.2 then translate into
$H_{-}\colon\theta_{i}<0,~i=1,\dots,K$
$H_{0}\colon\theta_{i}=0,~i=1,\dots,K$
$H_{+1}\colon\theta_{1}>0~\text{and}~\theta_{1}>\theta_{i},~i=2,\dots,K$
$\vdots$
$H_{+K}\colon\theta_{K}>0~\text{and}~\theta_{K}>\theta_{i},~i=1,\dots,K-1$
All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution. In analogy to the one treatment case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of corresponding hypothesis (e.g., for hypothesis $H_{+i}$, the space in $\mathbb{R}^{K}$ where the $i$th component is positive and larger than and all other components). Similarly, the prior hypothesis probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ may again be specified so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
Figure 4: Illustration of a spike-and-slab prior for a two-dimensional effect $\boldsymbol{\theta}=(\theta_{1},\theta_{2})^{\top}$. A point mass prior at $(0,0)^{\top}$ is assumed under $H_{0}$. A normal prior $\theta\sim\mathrm{N}((0,0)^{\top},\boldsymbol{\mathcal{T}})$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, with support truncated to the space of the corresponding hypothesis is assumed under $H_{-}$, $H_{+1}$, and $H_{+2}$. The correlation ensures that all treatments receive equal prior probability $\Pr(H_{-})=\Pr(H_{+1})=\Pr(H_{+2})=\{1-\Pr(H_{0})\}/3$.
Figure 4 illustrates such a spike-and-slab prior for the two treatment groups scenario ($K=2$). Specifying a prior covariance matrix $\boldsymbol{\mathcal{T}}$ with uniform correlation of 0.5 ensures that the prior probabilities of all hypotheses but $H_{0}$ are equal, which seems a sensible default. This can also be motivated by the fact that for normal outcomes with a shared control group and equal allocation, mean difference effect estimates are correlated by 0.5 due to the common control group for all treatments.
As in the two-group case (Section 3.1), the normal-normal conjugate framework allows us to derive marginal likelihoods in closed-form. The marginal likelihood under $H_{0}$ is
$p(\boldsymbol{\hat{\theta}}\mid H_{0})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{0},\boldsymbol{\Sigma}),$
while the marginal likelihood under $H_{-}$ is
$p(\boldsymbol{\hat{\theta}}\mid H_{-})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu}_{*},\boldsymbol{\mathcal{T}}_{*})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu},\boldsymbol{\mathcal{T}})}$
with $\mathrm{N}_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ and $\Phi_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ the density and cumulative distribution functions of the $K$-variate normal distribution with mean vector $\boldsymbol{m}$ and covariance matrix $\boldsymbol{V}$ evaluated at $\boldsymbol{x}$, and posterior mean $\boldsymbol{\mu}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}(\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\theta}}+\boldsymbol{\mathcal{T}}^{-1}\boldsymbol{\mu})$ and covariance $\boldsymbol{\mathcal{T}}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}$. Finally, the marginal likelihood under $H_{+i}$ is given by
$p(\boldsymbol{\hat{\theta}}\mid H_{+i})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu}_{*},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}_{*}\boldsymbol{A}_{i,K}^{\top})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}\boldsymbol{A}_{i,K}^{\top})}$
where $\boldsymbol{A}_{i,K}$ is a $K\times K$ contrast matrix that maps $\boldsymbol{\theta}$ to the space where the negative orthant corresponds to the space of hypothesis $H_{+i}$. For example, for $i=2$ and $K=3$, the matrix is
$\boldsymbol{A}_{2,3}=\begin{bmatrix}0&-1&0\\ 1&-1&0\\ 0&-1&1\\ \end{bmatrix}$
with the first row encoding the constraint of $\theta_{2}$ being positive, and the second and third rows encoding the constraints of $\theta_{2}$ being larger than $\theta_{1}$ and $\theta_{3}$, respectively. As expected, for $K=1$, these marginal likelihoods reduce to (7). Moreover, since they are all available in closed-form, Bayes factors, posterior probabilities, and randomization probabilities are also available in closed-form. Crucially, only the exact evaluation of multivariate normal densities and cumulative distribution functions is required, and both are efficiently implemented in statistical software (e.g., in the mvtnorm R package, Genz and Bretz,, 2009). This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve various complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
Figure 5: Evolution of Bayesian RAR probabilities for 3 treatments and a control group. In each step, one observation from each group is simulated from a normal distribution with a standard deviation of 1 and corresponding mean differences as indicated in the panel titles (treatment 1 is the most effective treatment). Randomization probabilities are computed using a normal spike-and-slab prior centered at the origin and with covariance matrix $\boldsymbol{\mathcal{T}}$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, and for different prior probabilities $\Pr(H_{0})$.
Figure 5 shows sequences of randomization probabilities computed from simulated normal data with $K=3$ treatment groups. We see again that assigning a prior probability $\Pr(H_{0})=1$ leads to static equal randomization at $\pi_{i}=1/(K+1)=25\%$, while assigning $\Pr(H_{0})=0$ (Thompson sampling) leads to the most variable randomization probabilities. Since data are simulated assuming that treatment 1 is the most effective, randomization probabilities based on $\Pr(H_{0})<1$ converge towards $\pi_{1}=100\%$ and toward $0\%$ for the remaining treatments. While convergence is the fastest for $\Pr(H_{0})=0$, this prior probability also accidentally produces rather high randomization probabilities for the control group at the start of the study, which is less pronounced for positive prior probabilities $\Pr(H_{0})>0$.
4 Point null Bayesian RAR for binary outcomes
We now consider the setting with binary outcomes, as such outcomes frequently occur in applications of Bayesian RAR (e.g., in clinical trials or A/B testing settings). Suppose that we observe data of the form $y=\{y_{C},y_{1},\dots,y_{K},n_{C},n_{1},\dots,n_{K}\}$ where $y_{i}$ denotes the number of successes out of $n_{i}$ trials in group $i\in\{C,1,\dots,K\}$ coming from a control group (index $C$) and $K$ treatment groups. All success counts are assumed to be binomially distributed with probabilities $\theta_{C},\theta_{1},\dots,\theta_{K}$, respectively, and higher values are assumed to indicate a higher benefit (e.g., a higher probability of disease recovery). The hypotheses of interest from Section 2.2 then translate into
$H_{-}\colon\theta_{C}>\theta_{i},~i\in\{1,\dots,K\}$
$H_{0}\colon\theta_{C}=\theta_{1}=\dots=\theta_{K}$
$H_{+1}\colon\theta_{1}>\theta_{i},~i\in\{C,2,\dots,K\}$
$\vdots$
$H_{+K}\colon\theta_{K}>\theta_{i},~i\in\{C,1,\dots,K-1\}$
In the approximate normal framework from Section 3, these hypotheses could be translated into hypotheses related to log odds ratios $\psi_{i}=\log\{\theta_{i}(1-\theta_{C})\}/\{(1-\theta_{i})\theta_{C}\}$, which can be estimated with logistic regression or other methods (we will provide a comparison with this approach below). However, such normal approximations can be inaccurate for small sample sizes and/or extreme probabilities close to zero/one. It is therefore preferable to compute randomization probabilities via the exact binomial distribution.
To compute posterior and randomization probabilities, it is necessary to compute the marginal likelihood of the observed data $y$ under the different hypotheses. The null hypothesis $H_{0}$ is no longer a simple hypothesis but requires specification of a prior for the common probability $\theta_{C}$. Assuming a beta prior $\theta_{C}\mid H_{0}\sim\mathrm{Beta}(a_{0},b_{0})$, the marginal likelihood of the observed data is
$\Pr(y\mid H_{0})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{0}+\sum_{j\in\{C,1,\dots,K\}}y_{i},b_{0}+\sum_{j\in\{C,1,\dots,K\}}n_{i}-\sum_{j\in\{C,1,\dots,K\}}y_{i})}{\mathrm{B}(a_{0},b_{0})}$
with $\mathrm{B}(a,b)=\int_{0}^{1}t^{a-1}(1-t)^{b-1}\mathrm{d}t$ the beta function. Under the remaining hypotheses, it is natural to assume independent beta priors $\theta_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ for $i\in\{C,1,\dots,K\}$, and truncate their support to the space of the corresponding hypothesis. This leads to the marginal likelihoods
$\Pr(y\mid H_{+i})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{j}+y_{j},b_{j}+n_{j}-y_{j})}{\mathrm{B}(a_{j},b_{j})}$
$\times\frac{Q_{i}(a_{C}+y_{C},a_{1}+y_{1},\dots,a_{K}+y_{K},b_{C}+n_{C}-y_{C},b_{1}+n_{1}-y_{1},\dots,b_{K}+n_{K}-y_{K})}{Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})}$ (9)
with
$Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ $=\Pr(\theta_{i}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}\mid a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$
$=\int_{0}^{1}p(\theta_{i}\mid a_{i},b_{i})\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}\Pr(\theta_{j}<\theta_{i}\mid a_{j},b_{j})\,\mathrm{d}\theta_{i}$
$=\int_{0}^{1}\frac{\theta_{i}^{a_{i}-1}(1-\theta_{i})^{b_{i}-1}}{\mathrm{B}(a_{i},b_{i})}\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}I_{\theta_{i}}(a_{j},b_{j})\,\mathrm{d}\theta_{i},$
where $I_{x}(a,b)=\{\int_{0}^{x}t^{a-1}(1-t)^{b-1}\,\mathrm{d}t\}/\mathrm{B}(a,b)$ is the regularized incomplete beta function, also known as the cumulative distribution function of the beta distribution. These can be efficiently computed with numerical integration and standard implementations of the regularized incomplete beta function (e.g., stats::pbeta in R). The marginal likelihood of the data under $H_{-}$ can similary be obtained from (9) with $i=C$.
For a specified prior probability of the null hypotheses $\Pr(H_{0})$, we may again distribute the remaining prior probability among the other hypotheses by
$\Pr(H_{+i})=\{1-\Pr(H_{0})\}\{1-Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})\}$
to ensure that the averaged prior is again the beta prior that was truncated in the first place. Plugging these marginal likelihoods and prior probabilities into equation (1) produces posterior probabilities, which in turn can be used for obtaining randomization probabilities. Similarly, ratios of marginal likelihoods can be taken to obtain Bayes factors for monitoring accumulating evidence.
Figure 6: Evolution of Bayesian RAR probabilities for 3 treatments and a control group. In each step, one observation from each group is simulated from a binomial distribution with probability as indicated in the panel titles (treatment 1 is the most effective). Randomization probabilities are computed assuming binomial likelihoods with independent uniform priors or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior with variances 11 and correlations $0.5$.
Figure 6 illustrates RAR probabilities for simulated binomial data with a control and $K=3$ treatment groups. Randomization probabilities were computed assuming exact binomial likelihoods with uniform priors for the probabilities or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior. We can see that the exact (solid lines) and approximate (dashed lines) probabilities are not too far from each other in most cases even though the true probabilities from which the data are simulated are rather small. The RAR probabilities with $\Pr(H_{0})=0.5$ remain considerably closer to 25% (equal randomization) compared to the RAR probabilities with $\Pr(H_{0})=0$ (Thompson sampling), which shows the largest variability. Since the true probability in treatment group 1 is the highest, all randomization probabilities converge towards $100\%$ for treatment 1, as expected, while they decrease to $0\%$ for the other treatments.
5 Reanalyzing the ECMO trial
The ECMO trial (Bartlett et al.,, 1985) investigated the efficacy of ECMO (extracorporeal membrane oxygenation) treatment in the critical care of newborns. It was the first prominent clinical trial to use a “randomized play-the-winner” (RPW) RAR design (Wei and Durham,, 1978). While some earlier non-randomized studies had shown a substantial treatment effect, a randomized study was required to confirm this. However, this presented an ethical dilemma, as the investigators were convinced that the risk of death would be much higher in the control group than in the ECMO group. To mitigate this, the RPW design was chosen.
The outcome of the trial was extreme: The first newborn was randomized to receive ECMO treatment and survived. The second newborn was randomized to receive the control treatment and died. The ten subsequently enrolled newborns were all randomized to ECMO, and all survived (Bartlett et al.,, 1985). The trial was then stopped for efficacy and its results published. However, the unusual outcome sparked intense debates about whether the trial was even a proper randomized clinical trial. Several subsequent trials were conducted, all re-establishing the effectiveness of ECMO, which is now a standard treatment (Bartlett,, 2024). In the following, we will reanalyze the ECMO data.
Figure 7: Evolution of Bayesian RAR randomization probabilities (top plot) and posterior probability of a beneficial ECMO treatment effect (bottom plot) for data from the ECMO trial.
The top plot in Figure 7 shows Bayesian RAR probabilities computed from the ECMO data sequence. The normal approximation and the exact binomial methods were used, as well as the RPW method that was originally used in the ECMO trial. A standard normal prior was assigned to the log odds ratio for the normal method, while uniform priors were assigned to the probabilities for the binomial method. For the normal method, log odds ratio estimates were computed with a Yates’ correction (i.e., adding a half to each cell) to avoid issues with zero cells. We can see that this correction induces a slight anomaly at the start of the study as the probabilities from the normal method slightly decrease after the first observation despite the first patient surviving the ECMO treatment. This does not happen for the exact method whose randomization probability increase after every additional patient, which only happens for the normal approximation after the second patient. Furthermore, the randomization probabilities differ notably between the normal and exact methods, even at the end of the study. This presumably happens because the influence of the differing priors remains relatively high after observing only 12 patients.
Comparing the randomization probabilities for different prior probabilities of the null hypothesis for each method, we can see that $\Pr(H_{0})=0$ (Thompson sampling) shows the most extreme randomization probabilities that rapidly increase to 100%, whereas $\Pr(H_{0})=1$ (equal randomization) leaves the probabilities completely static at 50%. In between, the randomization probabilities gradually shift from 50% to the Thompson sampling probabilities. The RPW probabilities (black squares) are relatively close to the probabilities from the exact method with $\Pr(H_{0})=0.5$ at the beginning of the study and become closer to the normal approximate method with $\Pr(H_{0})=0.75$ at later time points.
The bottom plot in Figure 7 shows the corresponding posterior probability of the ECMO treatment being effective. Depending on the prior probability $\Pr(H_{0})$, we can see that the posterior probability at the end of the trial may be very high (e.g., $\Pr(H_{+}\mid y)=0.99$ for $\Pr(H_{0})=0$), or only moderately favoring ECMO (e.g., $\Pr(H_{+}\mid y)=0.86$ for $\Pr(H_{0})=0.5$). From a Bayesian perspective, stopping the trial seems thus only a sensible decision if the prior probability of equal effects was low. Conversely, more evidence would have been needed if the prior probability had been higher, for example, if the prior probability were $\Pr(H_{0})=0.5$, representing a priori equipoise.
6 Simulation study
We conducted a simulation study to evaluate the performance of point null Bayesian RAR for different values of the prior probability of the null hypothesis $\Pr(H_{0})$, and compare it to Thompson sampling (potentially modified with burn-in periods, probability capping, power transformations) and equal randomization. Considered patient performance measures were the rate of successes, the rate of extreme randomization probabilities, and sample size imbalance. Additionally, bias and coverage were considered to evaluate the performance of rate difference point estimates and confidence intervals under RAR, while the type I error rate and power of the corresponding tests were used to quantify hypothesis testing performance under RAR. A binomial data-generating mechanism was used which was partially based on the simulation studies from Robertson et al., (2023), Thall and Wathen, (2007), and Wathen and Thall, (2017). Detailed description of the design and results of the simulation study following the structured ADEMP approach (Morris et al.,, 2019; Siepe et al.,, 2024) are provided in Appendix B. A supplemental website provides an interactively explorable results dashboard (https://samch93.github.io/brar/).
Across all simulations, we observed a trade-off between patient benefit and parameter estimation / hypothesis testing operating characteristics. Some methods performed better in terms of patient benefit but had worse bias, coverage, type I error rate, and power while others performed better in terms of bias, coverage, type I error rate, and power but had worse patient benefit. This trade-off is well described in the RAR literature (Hu and Rosenberger,, 2003; Zhang and Rosenberger,, 2005; Williamson and Villar,, 2019; Robertson et al.,, 2023).
The main result of the simulation study regarding the newly proposed method was that, under most conditions, point null Bayesian RAR with a high prior probability of the null hypothesis $\Pr(H_{0})=0.75$ showed similar operating characteristics to Thompson sampling with capped randomization probabilities at 10% and 90% and a power transformation with $c=i/(2n)$, where $i$ is the current sample size and $n$ is the maximum sample size. Both point null Bayesian RAR with a prior probability of $\Pr(H_{0})=0.75$ and modified Thompson sampling could mitigate some issues with ordinary Thompson sampling. For instance, they exhibited less negative sample size imbalance, biased parameter estimates, undercoverage, and inflated type I error rates. However, this improvement came at the cost of worse patient benefit performance compared to unmodified Thompson sampling. For instance, the mean success rate was lower, though still considerably better than equal randomization in most cases. Most importantly, the variability of randomization probababilitis was much reduced with $\Pr(H_{0})=0.75$, producing rarely negative imbalances (i.e., a large proportion of patients randomized to an inferior treatment). Setting a lower but positive prior probability than $\Pr(H_{0})=0.75$ produced operating characteristics comparable to those from less extreme modifications of Thompson sampling, such as a power transformation with $c=1/2$. In sum, the simulation study demonstrated that point null Bayesian RAR has comparable statistical properties to Thompson sampling with common ad hoc modifications.
7 Discussion
In this paper we have proposed a modification of standard Bayesian RAR (Thompson sampling) via recasting of the problem in a Bayesian hypothesis testing framework and the introduction of a point null hypothesis. While the plausibility of point null hypotheses is often a matter of philosophical debate (see e.g., Berger and Delampady,, 1987; Ly and Wagenmakers,, 2022), this method is useful in the RAR setting, as it can interpolate between equal randomization and Thompson sampling by changing the prior probability of the null hypothesis $\Pr(H_{0})$. This allows experimenters to balance patient benefit with classical operating characteristics, such as power, type I error rate, bias, and coverage. For large values of $\Pr(H_{0})$, we observed behaviors and operating characteristics similar to those obtained with ad hoc modifications of Thompson sampling, such as capping, burn-ins, and power transformations. Our method is implemented in the free and open source R package brar for binomial outcomes and for data summarized by approximately normal effect estimates. The latter makes the method applicable to many settings, for example, settings where treatment effects are estimated with regression analyses.
One advantage of our framework is that the randomization probabilities coherently correspond with the available statistical evidence (in the form of Bayes factors) and beliefs (in the form of posterior probabilities). In principle, both could also be used as decision-making tools instead of relying on frequentist test criteria. For instance, if the posterior probability of a treatment’s superiority is greater than 0.99, say, a study could be stopped. This also makes sense from the perspective that it is unnatural to randomize participants with extreme randomization probabilities that are associated with such high posterior probabilities.
Although we conducted a simulation study to understand the method’s basic behavior, more realistic and comprehensive evaluations are needed to understand its applicability in real-world conditions. For example, the method needs to be evaluated in combination with futility stopping (e.g., dropping of treatment arms which are shown to be ineffective at an interim analysis). Another issue to consider is how to select the prior probability of the null hypothesis. In our simulation study, we found that setting a value of $\Pr(H_{0})=0.75$ mitigated many of the issues with Thompson sampling. However, other choices could be considered, such as setting a higher value. Similar considerations apply to the prior distributions of the parameters, as we did not assess the effect of varying these on the operating characteristics. For instance, rather than setting the correlation of the multivariate normal prior to achieve equal randomization probabilities, it could be beneficial to set it so that the prior probability of the control being superior is always $\Pr(H_{-})=0.5$ and the remaining probability is distributed equally among the treatments. Future work may therefore investigate whether a more efficient RAR procedure can be obtained by specifying a certain prior distribution. Finally, in many clinical trial settings, RAR methods are not directly applicable because outcomes such as death may only be observed after long follow-up periods, by which time recruitment and randomized allocation will already have finished. An alternative could be to perform point null Bayesian RAR with an informative surrogate outcome that is sooner observed than the primary outcome (Gao et al.,, 2024).
Acknowledgments
We thank František Bartoš and Małgorzata Roos for valuable comments on drafts of the manuscript. The acknowledgment of these individuals does not imply their endorsement of the paper.
Conflict of interest
We declare no conflict of interest.
Software and data
Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/brar. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.17248628. We used the statistical programming language R version 4.5.0 (2025-04-11) for analyses (R Core Team,, 2024) along with the ggplot2 (Wickham,, 2016), dplyr (Wickham et al.,, 2023), SimDesign (Chalmers and Adkins,, 2020), mvtnorm (Genz and Bretz,, 2009), ggh4x (van den Brand,, 2024), ggpubr (Kassambara,, 2023), and knitr (Xie,, 2015) packages.
Appendix A The R package brar
Our R package can be installed by running remotes::install_github(repo = "SamCH93/brar", subdir = "package") in an R session (requires the remotes package, which is available on CRAN). The main functions of the package are brar_normal and brar_binomial, which implement the approximate normal method from Section 3 and the exact binomial method from Section 4. The following code chunk illustrates how the latter function can be used.
[…]
library(brar) # load package
## observed successes and trials in control and 3 treatment groups
y <- c(10, 9, 14, 13)
n <- c(20, 20, 22, 21)
## conduct exact point null Bayesian RAR
brar˙binomial(y = y, n = n,
              ## uniform prior for common probability under H0
              a0 = 1, b0 = 1,
              ## uniform priors for all probabilities
              a = c(1, 1, 1, 1), b = c(1, 1, 1, 1),
              ## prior probability of the null hypothesis
              pH0 = 0.5)
## DATA
##             Events Trials Proportion
## Control         10     20      0.500
## Treatment 1      9     20      0.450
## Treatment 2     14     22      0.636
## Treatment 3     13     21      0.619
## PRIOR PROBABILITIES
##    H-    H0   H+1   H+2   H+3
## 0.125 0.500 0.125 0.125 0.125
## BAYES FACTORS (BF_ij)
##         H-     H0   H+1    H+2   H+3
## H-   1.000 0.0341  2.16 0.1837 0.223
## H0  29.335 1.0000 63.45 5.3891 6.533
## H+1  0.462 0.0158  1.00 0.0849 0.103
## H+2  5.443 0.1856 11.77 1.0000 1.212
## H+3  4.490 0.1531  9.71 0.8249 1.000
## POSTERIOR PROBABILITIES
##      H-      H0     H+1     H+2     H+3
## 0.00777 0.91148 0.00359 0.04228 0.03488
## RANDOMIZATION PROBABILITIES
##     Control Treatment 1 Treatment 2 Treatment 3
##       0.236       0.231       0.270       0.263
[…]
Appendix B Simulation study
We now describe the design and results of our simulation study following the structured ADEMP approach (Morris et al.,, 2019; Siepe et al.,, 2024). Our simulation study was not preregistered as it constitutes early-phase methodological research where the properties of a new method are explored without the intention to give wide recommendations for practitioners (Heinze et al.,, 2023). A website with additional details and results is provided at https://samch93.github.io/brar/.
B.1 Aims
The aim of the simulation study is to evaluate the design characteristics of the newly proposed point null Bayesian RAR approach, and compare it to existing methods.
B.2 Data-generating mechanism
The data-generating mechanism was inspired by the simulation studies from Robertson et al., (2023), Thall and Wathen, (2007), and Wathen and Thall, (2017). In each repetition, a data set with $n$ binary outcomes is simulated through RAR: A patient $i$ is randomly allocated to the control group or one of the $K$ treatment groups based on randomization probabilities computed from the $1,\dots,i-1$ preceding outcomes. Depending on the allocation, an outcome is either simulated from a Bernoulli distribution with probability $\theta_{C}$ in the control group, $\theta_{1}$ in the first treatment group, or $\theta_{2}$ for the remaining treatment groups (in case $K>1$).
Parameters were chosen similar to the simulation study from Robertson et al., (2023). We vary the sample size $n\in\{200,654\}$ to represent low and high powered studies, the number of treatment groups $K\in\{1,2,3\}$, and probability in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$. The probability in the control group and the remaining groups is always fixed at $\theta_{C}=0.25$ and $\theta_{2}=\theta_{3}=0.3$, respectively. All these parameters are varied fully-factorially, leading to $2\times 3\times 3=18$ parameter conditions.
Since treatment allocation determines from which true probability an outcome is simulated, data generation is directly influenced by the RAR methods described below. These come with additional parameters that are, however, considered as method tuning parameters rather than true underlying parameters.
B.3 Estimands and other targets
The primary interest of this simulation study lies in assessing the patient benefit characteristics of different RAR methods. Additionally, the estimand of interest is the rate difference $\text{RD}_{1}=\theta_{1}-\theta_{C}$ and the target of interest is the null hypothesis of $\text{RD}_{1}=0$.
B.4 Methods
We consider the above described point null RAR methods. The prior probability of $H_{0}$ is a tuning parameter and controls the variability of the randomization probabilities. Setting $\Pr(H_{0})=1$ produces equal randomization, whereas $\Pr(H_{0})=0$ produces Thompson sampling. We consider values of $\Pr(H_{0})\in\{0,0.25,0.5,0.75,1\}$, as well as the normal approximation and exact binomial version of RAR. For approximate normal RAR, a normal prior with mean 0, variance 1, and in case of $K>1$ a correlation of 0.5, is considered. Independent uniform priors are assigned for binomial RAR. Log odds ratios along with their covariance are estimated with logistic regression and then used as inputs for the normal RAR method, while the exact method uses success counts and sample sizes only. In case a method fails to converge, equal randomization is applied as a back-up strategy, as this mimics what an experimenter might do in practice when a RAR method fails to converge (Pawel et al.,, 2025).
We also consider three modifications of these methods: In some conditions, a “burn-in” phase is carried out during which the first 50 patients are always randomized with equal probability $1/(K+1)$ to each group. For Thompson sampling ($\Pr(H_{0})=0$), we additionally consider conditions with power transformations of randomization probabilities, i.e., if $\pi_{k}$ is the randomization probability of group $k$, we take $\pi_{k}^{*}=\pi_{k}^{c}/\sum_{j\in\{C,1,\dots,K\}}\pi_{j}^{c}$. We consider $c=1/2$ and $c=i/(2n)$ with $i$ the current and $n$ the maximal sample size, which are two popular choices of the tuning parameter $c$ (Wathen and Thall,, 2017). Additionally, in some conditions “capping” is applied to Thompson sampling. That is, randomization probabilities outside the $10\%,90\%]$ interval are set to either 10% or 90%. After capping has been performed, randomization probabilities are re-normalized to sum to one (Wathen and Thall,, 2017; Lee and Lee,, 2021). This re-normalization is only performed for randomization probabilities greater than 10%, as these would otherwise be reduced again to probabilities less than 10%. In case, a re-normalized probability becomes less than 10%, it is also capped at 10% and second re-normalization performed. Finally, for equal randomization ($\Pr(H_{0})=1$), no burn-in, capping, or power transformation conditions are simulated as these manipulations have no effect.
B.5 Performance measures
Different performance measures were used. Patient benefit was quantified with:
  • The mean rate of successes per study
    $\overline{\text{RS}}=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\,\sum_{j=1}^{n}\frac{y_{ij}}{n}$
    where $y_{ij}$ denotes the 0/1 success indicator of patient $j$ in simulation $i$, $n$ is the sample size, and $n_{\text{sim}}$ is the number of simulation repetitions.
  • The mean rate of extreme randomization probabilities (less than 10% or greater than 90%)
    $\overline{\text{REP}}=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\sum_{j=1}^{n}\frac{\mathbb{1}(\text{any randomization probability at time }j~<10\%~\text{or}~>90\%)}{n}$
    with indicator function $\mathbb{1}(\cdot)$.
  • The proportion of simulations where the number of allocations to treatment 1 was at least 10% of the total sample size $n$ less than the average sample size in the remaining groups
    $\hat{S}_{0.1}=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\,\mathbb{1}\left(\frac{n-n_{1i}}{K}-n_{1i}>0.1n\right)$
    where $n_{1i}$ is the number of allocations to treatment group 1 in simulation $i$. For $K=1$, this reduces to the $\hat{S}_{0.1}$ sample size imbalance measure from Robertson et al., (2023), which in turn was inspired by the performance evaluation in the simulation study from Thall et al., (2015).
Parameter estimation and hypothesis testing performance was quantified with:
  • The empirical bias of the estimate of the rate difference $\text{RD}_{1}$
    $\mathrm{Bias}(\text{RD}_{1})=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\,\hat{\theta}_{1i}-\hat{\theta}_{0i}-\text{RD}_{1}$
    where $\hat{\theta}_{1i}$ and $\hat{\theta}_{Ci}$ are the maximum likelihood estimates of the probabilities $\theta_{1}$ and $\theta_{C}$ in simulation repetition $i$.
  • Empirical coverage of the $95\%$ Wald confidence intervals of $\text{RD}_{1}$
    $\mathrm{Coverage}(\text{RD}_{1})=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\,\mathbb{1}\left\{95\%~\text{CI includes}~\text{RD}_{1}~\text{in simulation}~i\right\}.$
  • Empirical rejection rate (type I error rate or power depending on the condition) related to the Wald test of the rate difference $\text{RD}_{1}$
    $\mathrm{RR}(\text{RD}_{1})=\frac{1}{n_{\text{sim}}}\sum_{i=1}^{n_{\text{sim}}}\,\mathbb{1}\left\{\text{Test rejects}~H_{0}\colon\text{RD}_{1}=0~\text{in simulation}~i\right\}.$
Each condition was simulated 10’000 times. This ensures a MCSE (Monte Carlo Standard Error) for Type I error rate, power, and coverage of at most 0.5%. MCSEs were calculated using the formulae from Siepe et al., (2024) and are provided for all measures in the following figures and the supplemental website.
B.6 Computational aspects
The simulation study was run on a server running Debian GNU/Linux 13 (trixie) and R version 4.5.0 (2025-04-11). The SimDesign R package was used to organize and run the simulation study (Chalmers and Adkins,, 2020). Our newly developed brar R package was used to perform Bayesian RAR. Code and data to reproduce this simulation study are available at https://github.com/SamCH93/brar.
B.7 Results
Convergence
Non-convergence happened only rarely for the normal approximation method due to logistic regression not converging at the start of the study when no events were observed for some groups. In this case, equal randomization was applied. The highest rate of such non-convergence was in a condition with $K=3$ treatment groups where 4.6% of the $n$ logistic regressions did not converge. No other forms of missingness were observed. Figures and tables with per-condition-method non-convergence rates are available at https://samch93.github.io/brar/.
Rate of successes and extreme randomization probabilities
The mean rate of successes is shown in Figure LABEL:fig:successrate. It was generally the highest for Thompson sampling and lowest for equal randomization. The normal approximation and exact method produced mostly similar rates, with the normal method sometimes showing slightly higher rates (e.g., for $K=3$ and $\text{RD}_{1}=0.2$). The Thompson sampling modifications generally reduced the mean success rate, with the greatest reduction achieved by combining capping and power transformation with $c=i/(2n)$. In conditions with small sample size these rates were similar as when the prior probability was $H_{0}=0.75$, while they were lower when the sample size was larger. This makes sense as for larger sample sizes, uncapped randomization probabilities are more likely to converge to extreme ones. This is also visible in Figure LABEL:fig:extreme where more extreme randomization probabilities are observed with increasing sample size and rate difference for all methods but the ones with capping.
Negative sample size imbalance
Figure LABEL:fig:imbalance shows negative sample size imbalance as quantified by the $\hat{S}_{0.1}$ metric. Negative imbalance was the greatest for Thompson sampling and reduced when modifications were applied. Similarly, increasing the prior probability of $H_{0}$ decreased negative imbalance, in some cases even below Thompson sampling with modifications.
Bias and coverage
Figures LABEL:fig:bias and LABEL:fig:coverage show empirical bias and coverage related to estimates of the rate difference between the first treatment group and the control group $\text{RD}_{1}$. In conditions where there was no difference ($\text{RD}_{1}=0$), all methods produced unbiased point estimates, though all methods but equal randomization also showed undercoverage. Bias occurred in conditions with non-zero rate differences, the bias being the greatest for Thompson sampling and decreasing to some extent when modifications were introduced or the prior probability of $H_{0}$ increased. Similarly, modifications or increasing prior probabilitities improved coverage, although they still remained suboptimal in most conditions. For large rate difference conditions, coverage was much better for the normal than the exact version of RAR. Similarly, bias was in some cases larger for the exact compared to the normal version.
Type I error rate and power
Figures LABEL:fig:t1e and LABEL:fig:power show empirical type I error rate and power associated with the Wald test of $\text{RD}_{1}=0$. We see that standard Thompson sampling shows an inflated type I error rate above the nominal 2.5% which is reduced to some extent by increasing either the prior probability of $H_{0}$ or applying modifications. In the same way, Thompson sampling exhibits reduced power compared to equal randomization, which is again alleviated by modifications or positive prior probability of the null hypotheses. In small sample sizes and large rate differences, power was slightly increased for the exact compared to the normal version of RAR, while for Thompson sampling the type I rate was slightly higher for the exact compared to the normal version.

### v2

@@
-Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization
+Stabilizing Thompson Sampling with Null Hypothesis Bayesian Response-Adaptive Randomization
 Abstract
-Response-adaptive randomization (RAR) methods use accumulated data to adapt randomization probabilities, aiming to increase the probability of allocating patients to effective treatments.
+Response-adaptive randomization (RAR) methods can be used to adapt randomization probabilities based on accumulating data, aiming to increase the probability of allocating patients to effective treatments.
 A popular RAR method is Thompson sampling, which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective.
-However, its high variability early in a trial can also increase the risk of assigning patients to inferior treatments.
-We propose a principled method based on Bayesian hypothesis testing to mitigate this issue.
-Specifically, we introduce a point null hypothesis that postulates equal effectiveness of treatments.
-This induces shrinkage toward equal randomization probabilities, with the degree of shrinkage controlled by the prior probability of the null hypothesis.
+However, its high variability can also increase the risk of assigning patients to inferior treatments and lead to inferential problems such as confidence interval undercoverage.
+We propose a principled method based on Bayesian hypothesis testing to address these issues: We introduce a null hypothesis postulating equal effectiveness of treatments.
+Bayesian model averaging then induces shrinkage toward equal randomization probabilities, with the degree of shrinkage controlled by the prior probability of the null hypothesis.
 Equal randomization and Thompson sampling arise as special cases when the prior probability is set to one or zero, respectively.
-Simulated and real-world examples illustrate that the proposed method balances highly variable Thompson sampling with static equal randomization.
-A simulation study demonstrates that the method can mitigate issues with ordinary Thompson sampling and has comparable statistical properties to Thompson sampling with common ad hoc modifications such as power transformation and probability capping.
-We implement the method in the open-source R package brar, enabling experimenters to easily perform point null Bayesian RAR and support more effective randomization of patients.
-Keywords: Adaptive trials, A/B testing, Bayes factor, Bayesian model averaging, sharp null hypothesis, spike-and-slab prior
+Simulated and real-world examples illustrate that the method balances highly variable Thompson sampling with static equal randomization.
+A simulation study demonstrates that the method can mitigate issues with Thompson sampling and has comparable statistical properties to Thompson sampling with common ad hoc modifications such as power transformation and probability capping.
+We implement the method in the free and open-source R package brar, enabling experimenters to easily perform null hypothesis Bayesian RAR and support more effective randomization of patients.
+Keywords: Adaptive trials, Bayes factor, Bayesian model averaging, outcome-adaptive randomization, spike-and-slab prior
 1 Introduction
-Response-adaptive randomization (RAR) methods, sometimes also called outcome-adaptive randomization methods, randomly allocate experimental units (e.g., patients or animals) to treatments in a manner that is informed by accumulating data (Thall and Wathen,, 2007; Berry et al.,, 2010; Grieve,, 2016; Robertson et al.,, 2023).
-A popular approach is Thompson sampling (Thompson,, 1933), which randomizes participants proportionally to the Bayesian posterior probability that each treatment is the most effective.
-Such RAR methods are attractive because they naturally balance gathering information on treatment effectiveness and assigning subjects to effective treatments.
-Despite its benefits, there are also various challenges with Thompson sampling.
-For example, the method can exhibit high variability, particularly in the early stages of a study when posterior uncertainty is high (Thall et al.,, 2015).
-This can lead to erratic allocation probabilities, ethical concerns (e.g., exposing participants to inferior treatments with high probability), and inferential challenges (e.g., reduction of statistical power or biased effect estimates).
-Consequently, there has been substantial interest in modifying Thompson sampling to make it more reliable.
-For example, Thall and Wathen, (2007) propose to use the randomization probability $\pi=p^{c}/\{p^{c}+(1-p)^{c}\}$ where $p$ is the posterior probability that the experimental treatment is more effective than the control treatment, and $c$ is an additional parameter that controls the variability of the method.
-Setting $c=1$ produces Thompson sampling, while $c<1$ reduces variability with $c=1/2$ being often recommended.
-Another approach to reduce extreme randomization probabilities is to cap them, for instance, setting them to 10% or 90% if a method assigns more extreme probabilities (Thall and Wathen,, 2007; Lee and Lee,, 2021).
-Finally, RAR methods are often combined with “burn-in” periods at the start of the study, during which units are randomized with equal probabilities to mitigate high variability (Thall and Wathen,, 2007; Wathen and Thall,, 2017; Robertson et al.,, 2023).
-While such ad hoc modifications can address some of the limitations of RAR methods, they conflict with the principles of coherent Bayesian learning.
-For example, a transformed or capped posterior probability does no longer correspond to an actual posterior probability, and it cannot be used as a genuine prior for future data.
-This raises the question of whether it is possible to devise a RAR method with desirable properties, such as reduced variability compared to Thompson sampling, that is coherent with Bayesian principles, and if so, how it relates to these ad hoc modifications.
-In this paper, we propose a novel RAR method that reduces variability in a coherent Bayesian manner, which we term “point null Bayesian RAR”.
-The idea is to consider a point null hypothesis postulating that treatments are equally effective.
-This is equivalent to using a “spike-and-slab” prior (Raftery et al.,, 1997), also known as “lump-and-smear” prior, which is a mixture of a point mass at equal effectiveness and a probability density elsewhere (Spiegelhalter et al.,, 2004, Section 5.5.4).
-The prior probability of the point null hypothesis determines the mixture weight.
-As we will show, setting this prior probability to zero produces equal randomization, whereas setting it to one produces Thompson sampling.
-The proposed method thus interpolates between equal randomization and Thompson sampling in a coherent Bayesian way.
+Response-adaptive randomization (RAR) methods randomly allocate patients to treatments based on accumulating data (Hu and Rosenberger,, 2006; Thall and Wathen,, 2007; Berry et al.,, 2010; Robertson et al.,, 2023).
+A popular approach is Thompson sampling (Thompson,, 1933), which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective.
+Such RAR methods are attractive as they balance gathering information on treatment effectiveness and assigning subjects to effective treatments.
+Despite its benefits, Thompson sampling can also introduce inferential problems such as inflated type I error rates or bias in effect estimation.
+Furthermore, although Thompson sampling allocates patients to the more effective treatment on average, it may also increase the probability of allocating patients to inferior treatments compared to equal randomization, particularly when treatment effects are small.
+This can undermine the method’s intended objective and raise ethical concerns (Thall et al.,, 2015; Robertson et al.,, 2023).
+Consequently, there has been interest in modifying Thompson sampling to address these issues.
+One popular modification uses the randomization probability $\pi=p^{c}/\{p^{c}+(1-p)^{c}\}$, where $p$ is the posterior probability that the experimental treatment is more effective than the control, and $c$ is a tuning parameter.
+Setting $c=1$ produces Thompson sampling while $c<1$ reduces variability.
+Another approach is to cap randomization probabilities, for instance, setting them to 10% or 90% if the method assigns more extreme probabilities.
+Finally, “burn-in” periods at the start of the study are commonly used, during which patients are randomized with equal probabilities to mitigate high variability (Wathen and Thall,, 2017).
+Several simulation studies have shown that appropriately modified Thompson sampling can overcome the limitations of the ordinary formulation (e.g., Du et al.,, 2017; Viele et al.,, 2019, 2020; Lee and Lee,, 2021; Tang et al.,, 2025).
+However, such ad hoc modifications conflict with principles of coherent Bayesian learning.
+For example, a capped posterior no longer corresponds to an actual posterior and cannot serve as a genuine prior for future data.
+This raises the question of whether a principled Bayesian RAR method can mitigate Thompson sampling’s variability, and if so, how it relates to ad hoc modifications.
+Principled Bayesian alternatives to Thompson sampling have also been proposed outside clinical trials, particularly in the multi-armed bandit literature.
+For instance, RAR based on the Gittins index prioritizes treatments with the greatest expected reward (Gittins et al.,, 2011), or the Bayesian upper confidence bound method allocates treatments based on posterior quantiles and offers theoretical performance guarantees (Kaufmann et al.,, 2012).
+These methods are primarily used in online decision problems but rarely in clinical trials, where additional constraints such as type I error rate control are central (Villar et al.,, 2015).
+In this paper, we propose a novel RAR method that reduces variability in a coherent Bayesian manner, which we term “null hypothesis Bayesian RAR”.
+The idea is to introduce a null hypothesis postulating that treatments are equally effective.
+This is often equivalent to using a “spike-and-slab” prior for the treatment effect parameter, which is a mixture of a point mass at equal effectiveness and a probability density elsewhere (Spiegelhalter et al.,, 2004, Section 5.5.4).
+The prior probability of the null hypothesis determines the mixture weight.
+Setting it to zero produces equal randomization, whereas setting it to one produces Thompson sampling.
+The method thus interpolates between equal randomization and Thompson sampling in a coherent Bayesian way.
 As a by-product, posterior probabilities and Bayes factors are also obtained.
-These can be used to monitor evidence of the effectiveness of each treatment in a manner that is aligned with the randomization probabilities.
-In the following Section 2 we introduce the general idea of the method in more detail, followed by tailoring it to the setting of approximately normal effect estimates (Section 3) and binary outcomes (Section 4).
-In Section 5, we then illustrate the method on data from the ECMO trial (Bartlett et al.,, 1985), followed by evaluating its statistical properties in a simulation study (Section 6).
-The paper ends with concluding remarks on advantages, limitations, and opportunities for future research (Section 7).
-Appendix A illustrates our R package brar for performing point null Bayesian RAR.
-Appendix B provides further details on our simulation study.
-2 Point null Bayesian RAR
-We now explain the general idea of point null Bayesian RAR, without going into specifics such as data distribution or computation (these will follow in the subsequent sections).
-Throughout we will assume that we have observed data $y$ and we want to use these to randomize a future experimental unit.
-We start with the basic but important setting of one control and one treatment group, and extend it afterwards to multiple treatment groups.
-2.1 Two group comparisons
-In case there is a control group and only one treatment group, we consider the hypotheses:
+These can be used to monitor evidence of treatment effectiveness in a manner that is aligned with the randomization probabilities.
+In the following Section 2 we introduce the general idea of the method, followed by tailoring it to normal (Section 3) and binary outcomes (Section 4).
+Section 5 illustrates the method on data from the ECMO trial, followed by evaluating its properties in a simulation study (Section 6).
+Concluding remarks are given in Section 7.
+The supplement illustrates our R package brar, and provides further details on theory and the simulation study.
+2 Null hypothesis Bayesian RAR
+Assume that we have observed data $y$ and want to use these to randomize a future patient. Suppose there is one control and one treatment group, and consider the hypotheses:
 $H_{-}\colon\text{Treatment is less effective than control}$
 $H_{0}\colon\text{Treatment and control are equally effective}$
-$H_{+}\colon\text{Treatment is more effective than control}$
-How exactly these statements are translated into statistical hypotheses related to parameters depends on the type of data and model used, but often relates to an effect size parameter being less, equal, or greater than zero.
-There may also be situations where there is no control group but only two competing treatments.
-In this case, the method detailed here is still applicable but with the control group replaced by a reference group (the choice of the reference may be somewhat arbitrary).
-The Bayesian posterior probability of a hypothesis $H_{i}\in\{H_{-},H_{0},H_{+}\}$ can then be calculated by
-$\Pr(H_{i}\mid y)=\frac{p(y\mid H_{i})\Pr(H_{i})}{\sum\limits_{j\in\{-,0,+\}}p(y\mid H_{j})\Pr(H_{j})}=\left\{\sum_{j\in\{-,0,+\}}\mathrm{BF}_{ji}(y)\,\frac{\Pr(H_{j})}{\Pr(H_{i})}\right\}^{-1}$ (1)
-where $\Pr(H_{i})$ is the prior probability of hypothesis $H_{i}$,
-$p(y\mid H_{i})=\int_{\Theta}p(y\mid\theta)\,p(\theta\mid H_{i})\,\mathrm{d}\theta$
-is the marginal likelihood of the data $y$ under $H_{i}$ obtained from marginalizing the likelihood $p(y\mid\theta)$ with respect to the prior distribution $p(\theta\mid H_{i})$ assigned to the model parameters $\theta\in\Theta$ under $H_{i}$, and
-$\mathrm{BF}_{ji}(y)=\frac{\Pr(H_{j}\mid y)}{\Pr(H_{i}\mid y)}\,\bigg/\,\frac{\Pr(H_{j})}{\Pr(H_{i})}=\frac{p(y\mid H_{j})}{p(y\mid H_{i})}$ (2)
-is the Bayes factor contrasting $H_{j}$ to $H_{i}$ (Jeffreys,, 1939; Good,, 1958).
+$H_{+}\colon\text{Treatment is more effective than control.}$
+How exactly these are translated into statistical hypotheses about parameters depends on the data model, but often relates to an effect size parameter being less, equal, or greater than zero.
+If there is no control group, the method can still be used with the control replaced by a reference group.
+This hypothesis setup is similar to Cai et al., (2013) but with the addition of $H_{0}$, which introduces a different behavior of the resulting randomization probabilities.
+The Bayesian posterior probability of hypothesis $H_{i}\in\{H_{-},H_{0},H_{+}\}$ is then
+$\Pr(H_{i}\,|\,y)=\frac{p(y\,|\,H_{i})\Pr(H_{i})}{\sum\limits_{j\in\{-,0,+\}}p(y\,|\,H_{j})\Pr(H_{j})}=1\,\bigg/\,\bigg\{\sum_{j\in\{-,0,+\}}\mathrm{BF}_{ji}(y)\,\frac{\Pr(H_{j})}{\Pr(H_{i})}\bigg\}$ (1)
+where $\Pr(H_{i})$ is the prior probability of $H_{i}$ (with $\sum_{i\in\{-,0,+\}}\Pr(H_{i})=1$), $p(y\,|\,H_{i})=\int_{\Theta}p(y\mid\theta)\,p(\theta\,|\,H_{i})\,\mathrm{d}\theta$ is the marginal likelihood of the data $y$ under $H_{i}$ obtained from integrating the likelihood $p(y\,|\,\theta)$ with respect to the prior distribution $p(\theta\,|\,H_{i})$ of the model parameters $\theta\in\Theta$ under $H_{i}$, and
+$\mathrm{BF}_{ji}(y)=\frac{\Pr(H_{j}\,|\,y)}{\Pr(H_{i}\,|\,y)}\,\bigg/\,\frac{\Pr(H_{j})}{\Pr(H_{i})}=\frac{p(y\,|\,H_{j})}{p(y\,|\,H_{i})}$ (2)
+is the Bayes factor contrasting $H_{j}$ to $H_{i}$ (Kass and Raftery,, 1995).
 The Bayes factor (2) is the updating factor of the prior odds of $H_{j}$ to $H_{i}$ to the corresponding posterior odds (first equality), which is equivalent to the ratio of marginal likelihoods of the data under $H_{j}$ and $H_{i}$ (second equality).
-The posterior probabilities (1) can thus be computed from the marginal likelihoods of the data under each considered hypotheses along with their prior probabilities, or from the set of Bayes factors and prior hypothesis odds relative to some reference hypothesis (Kass and Raftery,, 1995).
+The posterior probabilities (1) can thus be computed from the marginal likelihoods of the data under each considered hypotheses along with their prior probabilities, or from the Bayes factors and prior odds relative to some reference hypothesis.
 Regardless in which way they are computed, the question is how to translate posterior probabilities into randomization probabilities.
-We propose to randomize a future unit to the treatment group with probability
-$\pi=\Pr(H_{+}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2},$ (3)
-while the probability to randomize to the control group is consequently
-$1-\pi=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2}.$ (4)
-Figure 1 shows such randomization probabilities for different combinations of posterior probabilities.
-We can see that the randomization probability $\pi$ shrinks towards 50% as the posterior probability of the null hypothesis $\Pr(H_{0}\mid y)$ increases.
-For example, if $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$, as indicated by the black asterisk in Figure 1, the probability to randomize to the treatment group is $\pi=0.2+0.5/2=45\%$.
+We propose to randomize a future patient to the treatment group with probability
+$\pi=\Pr(H_{+}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,2,$ (3)
+while the probability to randomize to the control group is consequently $1-\pi=\Pr(H_{-}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,2$.
+Figure 1 shows $\pi$ for different combinations of posterior probabilities.
+We can see that $\pi$ shrinks towards 50% as the posterior probability of the null hypothesis $\Pr(H_{0}\,|\,y)$ increases.
+For example, if $\Pr(H_{+}\,|\,y)=0.2$, $\Pr(H_{-}\,|\,y)=0.3$, and $\Pr(H_{0}\,|\,y)=0.5$, as indicated by the black asterisk, the randomization probability is $\pi=0.2+0.5/2=45\%$.
 Figure 1: Ternary plot where the color of each point depicts the treatment randomization probability $\pi$ for a particular combination of posterior probabilities of $H_{+}$ (bottom axis), $H_{-}$ (left axis), and $H_{0}$ (right axis).
-For example, the black asterisk denotes the combination where $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$ with corresponding randomization probability $\pi=45\%$.
+For example, the black asterisk denotes the combination where $\Pr(H_{+}\,|\,y)=0.2$, $\Pr(H_{-}\,|\,y)=0.3$, and $\Pr(H_{0}\,|\,y)=0.5$ with corresponding randomization probability $\pi=45\%$.
 This scheme can be motivated as Bayesian hypothesis averaged randomization probability where the hypothesis-specific treatment randomization probabilities are $0\%$, $50\%$, and $100\%$, under $H_{-}$, $H_{0}$, and $H_{+}$, respectively.
-That is, if we would know that $H_{+}$ or $H_{-}$ is true, we should assign the next patient to the treatment or control group, respectively, with probability one to maximize utility (patient benefit).
-On the other hand, if we knew that $H_{0}$ is true, then randomizing the assignment with $\pi=50\%$ seems sensible.
-Interestingly, the randomization scheme (3) reduces to Thompson sampling when the prior probability of $H_{0}$ is set to zero since then $\Pr(H_{0}\mid y)=0$ regardless of the data, and consequently $\pi=\Pr(H_{+}\mid y)$.
-However, the scheme induces shrinkage toward randomization probabilities of $\pi=50\%$ otherwise.
-In the most extreme case when $\Pr(H_{0})=1$, equal randomization is obtained as then $\Pr(H_{0}\mid y)=1$ regardless of the data, and consequently $\pi=50\%$.
+That is, if we knew that $H_{+}$ or $H_{-}$ is true, we should assign the next patient to the treatment or control group, respectively, to maximize patient benefit.
+In contrast, if we knew that $H_{0}$ is true, randomizing with $\pi=50\%$ seems sensible.
+The randomization scheme (3) reduces to Thompson sampling when the prior probability of $H_{0}$ is set to zero since then $\Pr(H_{0}\,|\,y)=0$ regardless of the data, and consequently $\pi=\Pr(H_{+}\,|\,y)$.
+However, the scheme induces shrinkage toward $\pi=50\%$ otherwise.
+In the most extreme case when $\Pr(H_{0})=1$, equal randomization is obtained as then $\Pr(H_{0}\,|\,y)=1$ regardless of the data, and consequently $\pi=50\%$.
 The scheme thus interpolates between Thompson sampling and equal randomization in a coherent Bayesian way.
@@
 […]
-$\vdots$
-[…]
-Posterior probabilities of each hypothesis can be computed from the marginal likelihoods and prior hypotheses probabilities with the summation in (1) extended to encompass all hypotheses (i.e., summing over $j\in\{-,0,+1,\dots,+K\}$).
-Similarly, they can be translated into randomization probabilities
-$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1}$ (5)
-with corresponding control randomization probability
-$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1},$ (6)
-which reduce to the randomization probabilities (3) and (4) for $K=1$.
-Also in the multi-treatment case, the randomization probabilities are shrunken towards equal randomization $\pi=1/(K+1)$ by introducing the null hypothesis $H_{0}$.
-Similarly, Thompson sampling and equal randomization are obtained by setting $\Pr(H_{0})=0$ and $\Pr(H_{0})=1$, respectively, since then the posteriors $\Pr(H_{0}\mid y)=0$ and $\Pr(H_{0}\mid y)=1$ are obtained for any observed data $y$.
-In this paper, we focus on schemes (5) and (6), which induce shrinkage toward equal randomization.
-However, it is important to note that there are alternatives to this approach.
-For example, it may be desired to shrink to different “baseline” randomization probabilities than equal randomization probabilities.
-This can be achieved by modifying the multiplicative factor of $\Pr(H_{0}\mid y)$ in (5) from $1/(K+1)$ to the desired baseline randomization probability.
-For example, if the goal is to minimize the standard errors of the treatment effect estimates by using $\sqrt{K}:1:\dots:1$ square-root allocation of the control to treatments (Dunnett,, 1955), we may use
-$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+\sqrt{K}}$
-which leads to the control randomization probability
-$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{\sqrt{K}}{K+\sqrt{K}}$
-These correctly shrink RAR probabilitites towards the Dunnett-type randomization probabilities, $\pi_{i}=1/(K+\sqrt{K})$ for treatment $i=1,\dots,K$, and $1-\sum_{i=1}^{K}\pi_{i}=\sqrt{K}/(K+\sqrt{K})$ for control.
-3 Point null Bayesian RAR under approximate normality
-We will expand on point null Bayesian RAR in the setting where the data are summarized by an asymptotically normally distributed effect estimate.
-This does not mean that the raw data (e.g., a vector of outcomes and a matrix of covariates), from which the estimate is computed, need to be normally distributed.
-For instance, the regression coefficients from generalized linear models estimated using maximum likelihood satisfy asymptotic normality even if the data are themselves not normally distributed.
-This is useful because it allows us to efficiently compute posterior and randomization probabilities without the need for simulation.
-Although this framework is widely applicable, improvements may be possible by considering the exact distribution of the data.
-This will be detailed for binary outcomes in Section 4.
-3.1 Two group comparisons
-Assume that the data are summarized by $y=\{\hat{\theta},\sigma\}$, where $\hat{\theta}$ is an effect estimate of the true treatment effect $\theta$ that quantifies the effect of a treatment over the control and $\sigma$ is the standard error of the estimate.
-For example, $\hat{\theta}$ could be an estimated (standardized) mean difference, log odds/rate/hazard ratio, risk difference, or regression coefficient.
-Suppose further that the estimate is (at least approximately) normally distributed, i.e., $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2})$.
-If the effect $\theta$ is oriented such that a positive effect indicates treatment benefit, the three hypotheses from Section 2.1 translate into:
-$H_{-}\colon\theta<0$ versus $H_{0}\colon\theta=0$ versus $H_{+}\colon\theta>0$
-The point null hypothesis $H_{0}$ is a simple hypothesis with no free parameters, or equivalently, a point (Dirac) prior at zero.
-The $H_{-}$ and $H_{+}$ hypotheses are composite hypotheses and require prior distributions for the effect $\theta$.
-A natural choice is a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ normal distribution whose support is truncated to the negative and positive side, respectively.
-This distribution may be specified based on prior knowledge (e.g., previous studies).
-In the absence of prior knowledge, it seems sensible to center the prior on zero ($\mu=0$) to represent clinical equipoise (Freedman,, 1987), as a zero-centered prior gives equal probability to harmful and beneficial effects.
-Averaging the prior over the three hypotheses leads to a spike-and-slab prior, as illustrated in Figure 2.
+for $i=1,\dots,K,$ which span all effect ordering possibilities.
+Posterior hypothesis probabilities can still be computed from (1) but summing over all hypotheses (i.e., $j\in\{-,0,+1,\dots,+K\}$).
+These can then be translated into treatment randomization probabilities
+$\pi_{i}=\Pr(H_{+i}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+1)$ (4)
+and control randomization probability $1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+1)$.
+Also in the multi-treatment case, the randomization probabilities shrink towards equal randomization $\pi_{i}=1/(K+1)$ by introducing the null hypothesis $H_{0}$.
+Thompson sampling and equal randomization are again obtained by setting $\Pr(H_{0})=0$ and $\Pr(H_{0})=1$, respectively, as this leads to $\Pr(H_{0}\,|\,y)=0$ and $\Pr(H_{0}\,|\,y)=1$ for any data $y$.
+In this paper, we focus on scheme (4), which induces shrinkage toward equal randomization.
+However, it is important to note that there are alternatives.
+It may be desired to shrink to different “baseline” randomization probabilities.
+This can be achieved by modifying the factor of $\Pr(H_{0}\,|\,y)$ in (4) from $1/(K+1)$ to the desired baseline randomization probability.
+For example, the aim may be to test all $K$ pairwise differences between the treatments and the common control using Dunnett’s test.
+A square root allocation rule (with $\sqrt{K}$ more patients in the control group than in each treatment group) is commonly used to maximize the power of the test (Kieser,, 2020, p.156).
+We may therefore use $\pi_{i}=\Pr(H_{+i}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+\sqrt{K})$.
+These correctly shrink RAR probabilitites towards Dunnett-type randomization probabilities, $\pi_{i}=1/(K+\sqrt{K})$ for treatment $i=1,\dots,K$, and $1-\sum_{i=1}^{K}\pi_{i}=\sqrt{K}/(K+\sqrt{K})$ for control.
+2.2 Asymptotic randomization probabilities
+It is of interest to understand how RAR probabilities behave as data accumulate.
+Under ordinary non-RAR based sampling, the posterior distribution is “consistent”, meaning that as more data are generated under a hypothesis the corresponding posterior hypothesis probability tends to one (Dawid,, 2011).
+In Section 3, we provide empirical support through simulation that posterior consistency also holds for the RAR setting.
+Thus if a treatment is superior and $\Pr(H_{0})<1$, the corresponding posterior and randomization probability tend to one.
+Conversely, if all treatments are equally effective and $\Pr(H_{0})>0$, the posterior probability of $H_{0}$ will approach one, and the randomization probabilities will hence tend to the specified baseline randomization probabilities.
+This is an important difference to Thompson sampling, whose randomization probabilities do not converge to the baseline randomization probability under $H_{0}$ but remain random (Deliu and Villar,, 2025).
+3 Null hypothesis Bayesian RAR under normality
+Suppose that the data are summarized by $y=\{\hat{\theta},\sigma\}$, where $\hat{\theta}$ is an estimate of the true treatment effect $\theta$, and $\sigma$ is the standard error of the estimate.
+This does not mean that the raw data, from which the estimate is computed, need to be normally distributed.
+For example, $\hat{\theta}$ could be an estimated (standardized) mean difference, log odds/risk/rate/hazard ratio, risk difference, or regression coefficient.
+The standard error is often of the form $\sigma=\lambda/\sqrt{n}$, where $n$ is the effective sample size and $\lambda$ the standard deviation of one effective observation.
+Assume that the estimate is (at least approximately) normally distributed around $\theta$ with its variance $\sigma^{2}$ assumed to be known, i.e., $\hat{\theta}\,|\,\theta\sim\mathrm{N}(\theta,\sigma^{2})$.
+If the effect $\theta$ is oriented such that a positive effect indicates treatment benefit, the hypotheses from Section 2 translate into $H_{-}\colon\theta<0$ vs.
+$H_{0}\colon\theta=0$ vs.
+$H_{+}\colon\theta>0$.
+The null hypothesis $H_{0}$ is a simple hypothesis with no free parameters, or equivalently, a point (Dirac) prior at zero.
+The $H_{-}$ and $H_{+}$ hypotheses are composite hypotheses and require prior distributions for $\theta$.
+A natural choice is a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ distribution whose support is truncated to the negative and positive side, respectively.
+In the absence of prior knowledge, it is sensible to center the prior on zero ($\mu=0$) to represent clinical equipoise (Freedman,, 1987), giving equal probability to harmful and beneficial effects.
+Since $H_{0}$ is a point hypothesis, the prior of $\theta$ under $H_{-}$ and $H_{+}$ must be proper ($\tau<\infty$), as otherwise the posterior probability of $H_{0}$ is 1 for any data $y$.
+The value of the prior standard deviation $\tau$ must therefore be chosen so that the prior is not too vague (Kass and Raftery,, 1995).
+The prior probability of the null hypothesis $\Pr(H_{0})$ represents the a priori plausibility of an absent effect, and also acts as a tuning parameter that controls the variability of RAR.
+An intuitive default is $\Pr(H_{0})=0.5$, representing equipoise between absent and present effects (Johnson,, 2013).
+For a given $\Pr(H_{0})$, it is then natural to set the prior probabilities of $H_{+}$ and $H_{-}$ to $\Pr(H_{+})=\{1-\Pr(H_{0})\}\times\Phi(\mu/\tau)$ and $\Pr(H_{-})=\{1-\Pr(H_{0})\}\times\Phi(-\mu/\tau)$, with $\Phi(\cdot)$ the standard normal CDF, so that to the prior distribution averaged over $H_{+}$ and $H_{-}$ is again the $\mathrm{N}(\mu,\tau^{2})$ distribution.
+For example, specifying $\Pr(H_{0})=0.5$ and a zero-centered prior, we obtain $\Pr(H_{+})=\Pr(H_{-})=0.5\times 0.5=0.25.$ Averaging the prior over all hypotheses leads then to a spike-and-slab prior, as illustrated in Figure 2.
 Figure 2: Illustration of spike-and-slab prior for the effect $\theta$.
@@
 These priors are averaged assuming prior hypothesis probabilities $\Pr(H_{0})=0.5$, $\Pr(H_{+})=0.25$, and $\Pr(H_{-})=0.25$.
-To compute posterior hypothesis probabilities, specification of prior hypothesis probabilities is required.
-The prior probability of the null hypothesis $\Pr(H_{0})$ represents the a priori plausibility of an absent effect, and also acts as a tuning parameter that controls the degree of variability of RAR.
-An intuitive default is $\Pr(H_{0})=0.5$ as it represents the equipoise position of equal probability of an absent effect relative to a present (either harmful or beneficial) effect (Johnson,, 2013).
-For a given $\Pr(H_{0})$, it is then natural to set the prior probabilities of the other two hypotheses to $\Pr(H_{+})=\{1-\Pr(H_{0})\}\times\Phi(\mu/\tau)$ and $\Pr(H_{-})=\{1-\Pr(H_{0})\}\times\Phi(-\mu/\tau)$ so that to the prior distribution averaged over $H_{-}$ and $H_{+}$ is again the $\mathrm{N}(\mu,\tau^{2})$ normal distribution that was truncated in the first place.
-For example, when setting $\Pr(H_{0})=0.5$ and specifying a zero-centered prior ($\mu=0$) as in Figure 2, we obtain $\Pr(H_{+})=\Pr(H_{-})=0.5\times 0.5=0.25.$
-With the prior densities and prior hypothesis probabilities specified, we can compute the marginal likelihood of the observed effect estimate under each hypothesis, and in turn obtain posterior probabilities (1).
-In the conjugate normal likelihood and prior framework, all of them can be straightforwardly derived in closed-form.
-Denoting by $\mathrm{N}(x\mid m,v)$ the normal density function with mean $m$ and variance $v$ evaluated at $x$, the marginal likelihoods are given by
-$p(\hat{\theta}\mid H_{-})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(-\mu/\tau)}$ (7a)
-$p(\hat{\theta}\mid H_{0})$ $=\mathrm{N}(\hat{\theta}\mid 0,\sigma^{2})$ (7b)
-$p(\hat{\theta}\mid H_{+})$ $=\mathrm{N}(\hat{\theta}\mid\mu,\sigma^{2}+\tau^{2})\times\frac{\Phi(\mu_{*}/\tau_{*})}{\Phi(\mu/\tau)}$ (7c)
-with posterior mean and variance
-$\mu_{*}=\frac{\hat{\theta}/\sigma^{2}+\mu/\tau^{2}}{1/\sigma^{2}+1/\tau^{2}}$ and $\tau_{*}^{2}=\frac{1}{1/\sigma^{2}+1/\tau^{2}}.$
-Taking ratios of marginal likelihoods produces the Bayes factors
+With the priors specified, we can compute the marginal likelihood of the observed effect estimate under each hypothesis and obtain posterior probabilities (1).
+In the normal likelihood and prior framework, all are available in closed-form.
+Denoting by $\mathrm{N}(x\,|\,m,v)$ the normal density function with mean $m$ and variance $v$ evaluated at $x$, the marginal likelihoods are
+$p(\hat{\theta}\,|\,H_{-})$ $=\mathrm{N}(\hat{\theta}\,|\,\mu,\sigma^{2}+\tau^{2})\times\Phi(-\mu_{*}/\tau_{*})\,/\,\Phi(-\mu/\tau)$
+$p(\hat{\theta}\,|\,H_{0})$ $=\mathrm{N}(\hat{\theta}\,|\,0,\sigma^{2})$
+$p(\hat{\theta}\,|\,H_{+})$ $=\mathrm{N}(\hat{\theta}\,|\,\mu,\sigma^{2}+\tau^{2})\times\Phi(\mu_{*}/\tau_{*})\,/\,\Phi(\mu/\tau)$
+with posterior mean and variance $\mu_{*}=(\hat{\theta}/\sigma^{2}+\mu/\tau^{2})\tau^{2}_{*}$ and $\tau_{*}^{2}=1/(1/\sigma^{2}+1/\tau^{2})$.
+Taking ratios of marginal likelihoods leads to the Bayes factors
 […]
 […]
-and the Bayes factors for other hypothesis comparisons can be obtained by transitivity and reciprocity, for example, $\mathrm{BF}_{-0}=\mathrm{BF}_{+0}\,/\,\mathrm{BF}_{+-}$.
-Posterior probabilities can now be obtained by plugging the Bayes factors and prior odds into (1), leading to
-[…]
-[…]
-[…]
-These Bayes factors and posterior probabilities can be monitored as data accumulate to see how the evidence for the hypotheses change.
-Moreover, they can be used to define the randomization probabilities via (3) leading to
-$\pi$ $=\left(1+\frac{\Phi(-\mu_{*}/\tau_{*})}{\Phi(\mu_{*}/\tau_{*})}+\frac{\Pr(H_{0})}{1-\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{\hat{\theta}^{2}}{\sigma^{2}}-\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}\right\}\right]\,\times\,\frac{\sqrt{1+\tau^{2}/\sigma^{2}}}{\Phi(\mu_{*}/\tau_{*})}\right)^{-1}$
-$\phantom{=}+\left(1+\frac{1-\Pr(H_{0})}{\Pr(H_{0})}\,\times\,\exp\left[-\frac{1}{2}\left\{\frac{(\hat{\theta}-\mu)^{2}}{\sigma^{2}+\tau^{2}}-\frac{\hat{\theta}^{2}}{\sigma^{2}}\right\}\right]\,\bigg/\,\sqrt{1+\frac{\tau^{2}}{\sigma^{2}}}\right)^{-1}\,\bigg/\,2.$ (8)
-As expected, the randomization probability (3.1) approaches equal randomization as the prior probability of $H_{0}$ increases to one (i.e., $\pi\to 50\%$ as $\Pr(H_{0})\nearrow 1$), while it approaches the ordinary Bayesian posterior tail probability of $\theta>0$ based on a normal prior $\theta\sim\mathrm{N}(\mu,\tau^{2})$ as the prior probability of $H_{0}$ decreases to zero (i.e., $\pi\to\Phi(\mu_{*}/\tau_{*})$ as $\Pr(H_{0})\searrow 0$).
-Figure 3: Evolution of Bayesian RAR probabilities for one treatment and control group.
-In each step one observation from the control and one from the treatment group are simulated from a normal distribution with a true standard deviation of 1 and assuming a true mean difference $\theta=0.25$.
-Randomization probabilities are then computed assuming a normal spike-and-slab prior centered at zero with standard deviation $\tau=1$, and different prior probabilities $\Pr(H_{0})$.
-Figure 3 shows sequences of randomization probabilities computed from simulated normal data.
-To enable comparison of randomization probabilities across different prior probabilities $\Pr(H_{0})$, the data were not simulated under RAR but by simulating an additional observation from the treatment and control groups at each step.
-We can see that the probability to randomize to the treatment group based on $\Pr(H_{0})=1$ remains static at $50\%$ (yellow line).
-In contrast, the randomization probabilities based on $\Pr(H_{0})<1$ tend towards $100\%$ as more data accumulate, as expected, given that the data were simulated under a beneficial treatment effect.
-Moreover, a clear ordering is visible: Probabilities under $\Pr(H_{0})=0$, which corresponds to Thompson sampling, are the most extreme and may even go strongly in the “wrong” direction.
-For example, the probability to randomize to control is higher than 75% at some point during the earlier stages of the study.
-In contrast, probabilities based on $0<\Pr(H_{0})<1$ show the same qualitative behavior but are less extreme.
-Setting a higher prior probability $\Pr(H_{0})$ thus reduces the variability of randomization probabilities but also makes convergence to a probability of 100% for the more effective treatment slower.
+and the Bayes factors for other hypothesis comparisons can be obtained by transitivity and reciprocity (e.g., $\mathrm{BF}_{-0}=\mathrm{BF}_{+0}\,/\,\mathrm{BF}_{+-}$).
+Posterior probabilities can now be obtained by plugging the Bayes factors and prior odds into (1).
+Bayes factors and posterior probabilities can be monitored as data accumulate to see how the evidence for the hypotheses changes.
+Moreover, they can be plugged into (3) to obtain the randomization probability
+$\pi$ $=\frac{\overbrace{\Phi\left(\mu_{*}/\tau_{*}\right)}^{\text{Thompson Sampling}}}{1+\frac{\Pr(H_{0})}{1-\Pr(H_{0})}\times\mathrm{BF}_{01}(\hat{\theta})}+\frac{\overbrace{1/2}^{\text{Equal Randomization}}}{1+\frac{1-\Pr(H_{0})}{\Pr(H_{0})}\,\big/\,\mathrm{BF}_{01}(\hat{\theta})}$ (5)
+where $\mathrm{BF}_{01}(\hat{\theta})=\exp[-\{(\hat{\theta}/\sigma)^{2}-(\hat{\theta}-\mu)^{2}/(\sigma^{2}+\tau^{2})\}/2]\sqrt{1+\tau^{2}/\sigma^{2}}$ is the Bayes factor contrasting $H_{0}$ to the alternative $H_{1}=H_{-}\cup H_{+}$ with prior $\theta\,|\,H_{1}\sim\mathrm{N}(\mu,\tau^{2})$.
+As expected, the randomization probability (5) shrinks towards equal randomization as the prior probability of $H_{0}$ increases (i.e., $\pi\to 50\%$ as $\Pr(H_{0})\nearrow 1$), whereas it approaches the ordinary Bayesian posterior tail probability of $\theta>0$ based on a $\theta\sim\mathrm{N}(\mu,\tau^{2})$ prior as the prior probability of $H_{0}$ decreases (i.e., $\pi\to 1-\Phi\{(0-\mu_{*})/\tau_{*}\}=\Phi(\mu_{*}/\tau_{*})$ as $\Pr(H_{0})\searrow 0$).
+Similarly, the randomization probability shrinks towards 50% as the data provide more evidence for $H_{0}$ ($\pi\to 50\%$ as $\mathrm{BF}_{01}\nearrow\infty$), whereas it approaches the ordinary posterior tail probability as the evidence for $H_{1}$ increases ($\pi\to\Phi(\mu_{*}/\tau_{*})$ as $\mathrm{BF}_{01}\searrow 0$).
+In contrast to ad hoc modifications of Thompson sampling, shrinkage of the posterior tail probability thus depends directly on the accumulated data via the Bayes factor.
+Figure 3: Evolution of Bayesian RAR probabilities for three simulated data sequences (a).
+In each step, one patient is allocated to the treatment or control group using Thompson sampling assuming a normal prior centered at zero with standard deviation $\tau=1$.
+Patient outcomes are then simulated from a normal distribution with a true standard deviation of 1 and assuming a true mean difference $\theta$ between treatment and control group as indicated in the plot panels.
+Randomization probabilities under the spike-and-slab prior ($\Pr(H_{0})>0$) are computed from the same data sequence to enable comparability.
+Plot (b) shows the distribution of randomization probabilities based on 1’000 simulations under Bayesian RAR for two different prior standard deviations and $\Pr(H_{0})$.
+Figure 3a shows sequences of randomization probabilities computed from normal data simulated under different true mean differences between treatment and control group $\theta$ as indicated in the panels.
+For each $\theta$, one sequence was simulated under RAR using Thompson sampling ($\Pr(H_{0})=0$).
+RAR probabilities for other values of $\Pr(H_{0})$ were computed from the same sequences to base comparison on the same data.
+Across all $\theta$, the probability to randomize to the treatment group based on $\Pr(H_{0})=1$ remains static at $50\%$ (yellow).
+In contrast, for a beneficial treatment effect ($\theta>0$; left plot), the randomization probabilities based on $\Pr(H_{0})<1$ tend towards $100\%$ as more data accumulate, whereas for a harmful treatment effect ($\theta>0$; right plot), they tend towards 0%.
+Moreover, a clear ordering is visible: Probabilities under $\Pr(H_{0})=0$ (Thompson sampling) are the most extreme both in the “correct” and “wrong” direction.
+Probabilities based on $0<\Pr(H_{0})<1$ show the same qualitative behavior but are less extreme.
+Setting a higher $\Pr(H_{0})$ thus reduces the variability of randomization probabilities but also slows down convergence to 100% or 0%.
+Finally, when there is no treatment effect ($\theta=0$; middle plot), the probabilities based on $0<\Pr(H_{0})<1$ stay relatively close to 50%, whereas the Thompson sampling probabilities tend to wander around.
+Figure 3b shows the distribution of randomization probabilities obtained from 1’000 simulations of the scenarios from Figure 3a and for two prior standard deviations $\tau$.
+When the true means differ from zero ($\theta=\pm 0.25$; first and third row), we see that the randomization probabilities for both $\Pr(H_{0})=0$ (Thompson sampling) and $\Pr(H_{0})=0.5$ correctly converge to 100% and 0%, respectively, although the convergence of $\Pr(H_{0})=0.5$ is slower.
+In contrast, when there is no difference ($\theta=0$; second row), only the randomization probabilities based on $\Pr(H_{0})=0.5$ converge to 50%, whereas the Thompson sampling randomization probabilities remain roughly uniformly distributed for all sample sizes.
+The qualitative behavior of the randomization probabilities is the same for both values of the prior standard deviation $\tau$.
+However, convergence to 100% and 0% is slightly faster for smaller $\tau$ (blue), whereas convergence to 50% is faster for larger $\tau$ (orange).
 3.1 More than two groups
-Suppose now there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to the control.
-A natural generalization is to stack the estimates into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume an approximate $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\mid\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the covariance matrix of $\boldsymbol{\hat{\theta}}$.
-For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ the corresponding covariance matrix.
-The hypotheses from Section 2.2 then translate into
-$H_{-}\colon\theta_{i}<0,~i=1,\dots,K$
-$H_{0}\colon\theta_{i}=0,~i=1,\dots,K$
-$H_{+1}\colon\theta_{1}>0~\text{and}~\theta_{1}>\theta_{i},~i=2,\dots,K$
-$\vdots$
-$H_{+K}\colon\theta_{K}>0~\text{and}~\theta_{K}>\theta_{i},~i=1,\dots,K-1$
+Suppose now there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to control.
+A natural generalization is to stack them into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume a $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\,|\,\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the (assumed to be known) covariance matrix of $\boldsymbol{\hat{\theta}}$.
+For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ its estimated covariance matrix.
+The hypotheses from Section 2.1 then translate into $H_{-}\colon\max\{\theta_{1},\dots,\theta_{K}\}<0$ vs.
+$H_{0}\colon\theta_{1}=\dots=\theta_{K}=0$ vs.
+$H_{+i}\colon\theta_{i}=\max\{\theta_{1},\dots,\theta_{K}\}>0$ for $i=1,\dots,K$.
 All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution.
-In analogy to the one treatment case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of corresponding hypothesis (e.g., for hypothesis $H_{+i}$, the space in $\mathbb{R}^{K}$ where the $i$th component is positive and larger than and all other components).
-Similarly, the prior hypothesis probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ may again be specified so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
+In analogy to the $K=1$ case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of the corresponding hypothesis.
+Similarly, a sensible default is to center the prior on $\boldsymbol{0}$ to encode clinical equipoise, and specify prior probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
 Figure 4: Illustration of a spike-and-slab prior for a two-dimensional effect $\boldsymbol{\theta}=(\theta_{1},\theta_{2})^{\top}$.
@@
 A normal prior $\theta\sim\mathrm{N}((0,0)^{\top},\boldsymbol{\mathcal{T}})$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, with support truncated to the space of the corresponding hypothesis is assumed under $H_{-}$, $H_{+1}$, and $H_{+2}$.
-The correlation ensures that all treatments receive equal prior probability $\Pr(H_{-})=\Pr(H_{+1})=\Pr(H_{+2})=\{1-\Pr(H_{0})\}/3$.
-Figure 4 illustrates such a spike-and-slab prior for the two treatment groups scenario ($K=2$).
-Specifying a prior covariance matrix $\boldsymbol{\mathcal{T}}$ with uniform correlation of 0.5 ensures that the prior probabilities of all hypotheses but $H_{0}$ are equal, which seems a sensible default.
-This can also be motivated by the fact that for normal outcomes with a shared control group and equal allocation, mean difference effect estimates are correlated by 0.5 due to the common control group for all treatments.
-As in the two-group case (Section 3.1), the normal-normal conjugate framework allows us to derive marginal likelihoods in closed-form. The marginal likelihood under $H_{0}$ is
-$p(\boldsymbol{\hat{\theta}}\mid H_{0})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{0},\boldsymbol{\Sigma}),$
-while the marginal likelihood under $H_{-}$ is
-$p(\boldsymbol{\hat{\theta}}\mid H_{-})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu}_{*},\boldsymbol{\mathcal{T}}_{*})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu},\boldsymbol{\mathcal{T}})}$
-with $\mathrm{N}_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ and $\Phi_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ the density and cumulative distribution functions of the $K$-variate normal distribution with mean vector $\boldsymbol{m}$ and covariance matrix $\boldsymbol{V}$ evaluated at $\boldsymbol{x}$, and posterior mean $\boldsymbol{\mu}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}(\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\theta}}+\boldsymbol{\mathcal{T}}^{-1}\boldsymbol{\mu})$ and covariance $\boldsymbol{\mathcal{T}}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}$.
-Finally, the marginal likelihood under $H_{+i}$ is given by
-$p(\boldsymbol{\hat{\theta}}\mid H_{+i})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu}_{*},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}_{*}\boldsymbol{A}_{i,K}^{\top})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}\boldsymbol{A}_{i,K}^{\top})}$
-where $\boldsymbol{A}_{i,K}$ is a $K\times K$ contrast matrix that maps $\boldsymbol{\theta}$ to the space where the negative orthant corresponds to the space of hypothesis $H_{+i}$.
-For example, for $i=2$ and $K=3$, the matrix is
-$\boldsymbol{A}_{2,3}=\begin{bmatrix}0&-1&0\\ 1&-1&0\\ 0&-1&1\\ \end{bmatrix}$
-with the first row encoding the constraint of $\theta_{2}$ being positive, and the second and third rows encoding the constraints of $\theta_{2}$ being larger than $\theta_{1}$ and $\theta_{3}$, respectively.
-As expected, for $K=1$, these marginal likelihoods reduce to (7).
-Moreover, since they are all available in closed-form, Bayes factors, posterior probabilities, and randomization probabilities are also available in closed-form.
-Crucially, only the exact evaluation of multivariate normal densities and cumulative distribution functions is required, and both are efficiently implemented in statistical software (e.g., in the mvtnorm R package, Genz and Bretz,, 2009).
-This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve various complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
-Figure 5: Evolution of Bayesian RAR probabilities for 3 treatments and a control group.
-In each step, one observation from each group is simulated from a normal distribution with a standard deviation of 1 and corresponding mean differences as indicated in the panel titles (treatment 1 is the most effective treatment).
-Randomization probabilities are computed using a normal spike-and-slab prior centered at the origin and with covariance matrix $\boldsymbol{\mathcal{T}}$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, and for different prior probabilities $\Pr(H_{0})$.
-Figure 5 shows sequences of randomization probabilities computed from simulated normal data with $K=3$ treatment groups.
-We see again that assigning a prior probability $\Pr(H_{0})=1$ leads to static equal randomization at $\pi_{i}=1/(K+1)=25\%$, while assigning $\Pr(H_{0})=0$ (Thompson sampling) leads to the most variable randomization probabilities.
-Since data are simulated assuming that treatment 1 is the most effective, randomization probabilities based on $\Pr(H_{0})<1$ converge towards $\pi_{1}=100\%$ and toward $0\%$ for the remaining treatments.
-While convergence is the fastest for $\Pr(H_{0})=0$, this prior probability also accidentally produces rather high randomization probabilities for the control group at the start of the study, which is less pronounced for positive prior probabilities $\Pr(H_{0})>0$.
-4 Point null Bayesian RAR for binary outcomes
-We now consider the setting with binary outcomes, as such outcomes frequently occur in applications of Bayesian RAR (e.g., in clinical trials or A/B testing settings).
-Suppose that we observe data of the form $y=\{y_{C},y_{1},\dots,y_{K},n_{C},n_{1},\dots,n_{K}\}$ where $y_{i}$ denotes the number of successes out of $n_{i}$ trials in group $i\in\{C,1,\dots,K\}$ coming from a control group (index $C$) and $K$ treatment groups.
-All success counts are assumed to be binomially distributed with probabilities $\theta_{C},\theta_{1},\dots,\theta_{K}$, respectively, and higher values are assumed to indicate a higher benefit (e.g., a higher probability of disease recovery).
-The hypotheses of interest from Section 2.2 then translate into
-$H_{-}\colon\theta_{C}>\theta_{i},~i\in\{1,\dots,K\}$
-$H_{0}\colon\theta_{C}=\theta_{1}=\dots=\theta_{K}$
-$H_{+1}\colon\theta_{1}>\theta_{i},~i\in\{C,2,\dots,K\}$
-$\vdots$
-$H_{+K}\colon\theta_{K}>\theta_{i},~i\in\{C,1,\dots,K-1\}$
-In the approximate normal framework from Section 3, these hypotheses could be translated into hypotheses related to log odds ratios $\psi_{i}=\log\{\theta_{i}(1-\theta_{C})\}/\{(1-\theta_{i})\theta_{C}\}$, which can be estimated with logistic regression or other methods (we will provide a comparison with this approach below).
-However, such normal approximations can be inaccurate for small sample sizes and/or extreme probabilities close to zero/one.
-It is therefore preferable to compute randomization probabilities via the exact binomial distribution.
-To compute posterior and randomization probabilities, it is necessary to compute the marginal likelihood of the observed data $y$ under the different hypotheses.
-The null hypothesis $H_{0}$ is no longer a simple hypothesis but requires specification of a prior for the common probability $\theta_{C}$.
-Assuming a beta prior $\theta_{C}\mid H_{0}\sim\mathrm{Beta}(a_{0},b_{0})$, the marginal likelihood of the observed data is
-$\Pr(y\mid H_{0})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{0}+\sum_{j\in\{C,1,\dots,K\}}y_{i},b_{0}+\sum_{j\in\{C,1,\dots,K\}}n_{i}-\sum_{j\in\{C,1,\dots,K\}}y_{i})}{\mathrm{B}(a_{0},b_{0})}$
-with $\mathrm{B}(a,b)=\int_{0}^{1}t^{a-1}(1-t)^{b-1}\mathrm{d}t$ the beta function.
-Under the remaining hypotheses, it is natural to assume independent beta priors $\theta_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ for $i\in\{C,1,\dots,K\}$, and truncate their support to the space of the corresponding hypothesis.
-This leads to the marginal likelihoods
-$\Pr(y\mid H_{+i})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{j}+y_{j},b_{j}+n_{j}-y_{j})}{\mathrm{B}(a_{j},b_{j})}$
-$\times\frac{Q_{i}(a_{C}+y_{C},a_{1}+y_{1},\dots,a_{K}+y_{K},b_{C}+n_{C}-y_{C},b_{1}+n_{1}-y_{1},\dots,b_{K}+n_{K}-y_{K})}{Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})}$ (9)
-with
-$Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ $=\Pr(\theta_{i}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}\mid a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$
-$=\int_{0}^{1}p(\theta_{i}\mid a_{i},b_{i})\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}\Pr(\theta_{j}<\theta_{i}\mid a_{j},b_{j})\,\mathrm{d}\theta_{i}$
-$=\int_{0}^{1}\frac{\theta_{i}^{a_{i}-1}(1-\theta_{i})^{b_{i}-1}}{\mathrm{B}(a_{i},b_{i})}\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}I_{\theta_{i}}(a_{j},b_{j})\,\mathrm{d}\theta_{i},$
-where $I_{x}(a,b)=\{\int_{0}^{x}t^{a-1}(1-t)^{b-1}\,\mathrm{d}t\}/\mathrm{B}(a,b)$ is the regularized incomplete beta function, also known as the cumulative distribution function of the beta distribution.
-These can be efficiently computed with numerical integration and standard implementations of the regularized incomplete beta function (e.g., stats::pbeta in R).
-The marginal likelihood of the data under $H_{-}$ can similary be obtained from (9) with $i=C$.
-For a specified prior probability of the null hypotheses $\Pr(H_{0})$, we may again distribute the remaining prior probability among the other hypotheses by
-$\Pr(H_{+i})=\{1-\Pr(H_{0})\}\{1-Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})\}$
-to ensure that the averaged prior is again the beta prior that was truncated in the first place.
-Plugging these marginal likelihoods and prior probabilities into equation (1) produces posterior probabilities, which in turn can be used for obtaining randomization probabilities.
-Similarly, ratios of marginal likelihoods can be taken to obtain Bayes factors for monitoring accumulating evidence.
-Figure 6: Evolution of Bayesian RAR probabilities for 3 treatments and a control group.
-In each step, one observation from each group is simulated from a binomial distribution with probability as indicated in the panel titles (treatment 1 is the most effective).
-Randomization probabilities are computed assuming binomial likelihoods with independent uniform priors or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior with variances 11 and correlations $0.5$.
-Figure 6 illustrates RAR probabilities for simulated binomial data with a control and $K=3$ treatment groups.
-Randomization probabilities were computed assuming exact binomial likelihoods with uniform priors for the probabilities or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior.
-We can see that the exact (solid lines) and approximate (dashed lines) probabilities are not too far from each other in most cases even though the true probabilities from which the data are simulated are rather small.
-The RAR probabilities with $\Pr(H_{0})=0.5$ remain considerably closer to 25% (equal randomization) compared to the RAR probabilities with $\Pr(H_{0})=0$ (Thompson sampling), which shows the largest variability.
-Since the true probability in treatment group 1 is the highest, all randomization probabilities converge towards $100\%$ for treatment 1, as expected, while they decrease to $0\%$ for the other treatments.
+The correlation of 0.5 ensures that all treatments receive equal prior probability $\Pr(H_{-})=\Pr(H_{+1})=\Pr(H_{+2})=\{1-\Pr(H_{0})\}/3$.
+Figure 4 illustrates a spike-and-slab prior for two treatment groups ($K=2$).
+The prior covariance matrix $\boldsymbol{\mathcal{T}}$ is set to a uniform correlation of 0.5 which ensures that the prior probabilities of all hypotheses but $H_{0}$ are equal.
+Note that a zero-centered prior with uniform correlation of 0.5 ensures equal treatment prior probabilities for any number of treatment groups $K$, and thus is a sensible default in the absence of prior knowledege.
+Also for $K>1$, marginal likelihoods, Bayes factors, posterior probabilities, and randomization probabilities are available in closed-form (shown in the supplement).
+Crucially, only the evaluation of multivariate normal densities and CDFs is required.
+This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve additional complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
+4 Null hypothesis Bayesian RAR for binary outcomes
+Suppose now that we observe binary data of the form $y=\{y_{C},y_{1},\dots,y_{K},n_{C},n_{1},\dots,n_{K}\}$ where $y_{i}$ denotes the number of successes out of $n_{i}$ trials in group $i\in\{C,1,\dots,K\}$ coming from a control group ($C$) and $K$ treatment groups.
+All success counts are assumed to be binomially distributed with probabilities $\theta_{C},\theta_{1},\dots,\theta_{K}$, respectively, with higher values indicating greater benefit.
+The hypotheses from Section 2.1 then translate into $H_{-}\colon\theta_{C}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}$ vs.
+$H_{0}\colon\theta_{C}=\theta_{1}=\dots=\theta_{K}$ vs.
+$H_{+i}\colon\theta_{i}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}$ for $i=1,\dots,K$.
+In the normal framework from Section 3, these could be translated into hypotheses related to log odds ratios, which can be estimated with logistic regression or other methods.
+However, normal approximations can be inaccurate for small sample sizes or extreme probabilities, so exact binomial computation is preferable.
+The null hypothesis $H_{0}$ requires specification of a prior for the common probability $\theta_{C}$.
+Assuming a beta prior $\theta_{C}\,|\,H_{0}\sim\mathrm{Beta}(a_{0},b_{0})$, the marginal likelihood under $H_{0}$ is
+$\Pr(y\,|\,H_{0})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{0}+\sum_{i\in\{C,1,\dots,K\}}y_{i},b_{0}+\sum_{i\in\{C,1,\dots,K\}}n_{i}-y_{i})}{\mathrm{B}(a_{0},b_{0})}$
+with beta function $\mathrm{B}(\cdot,\cdot)$.
+Under the other hypotheses, it is natural to assign independent beta priors $\theta_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ for $i\in\{C,1,\dots,K\}$, and truncate their support to the space of the corresponding hypothesis.
+This leads to the marginal likelihood of the data
+$\Pr(y\,|\,H_{i})$ $=\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{j}+y_{j},b_{j}+n_{j}-y_{j})}{\mathrm{B}(a_{j},b_{j})}$
+$\times\frac{Q_{i}(a_{C}+y_{C},a_{1}+y_{1},\dots,a_{K}+y_{K},b_{C}+n_{C}-y_{C},b_{1}+n_{1}-y_{1},\dots,b_{K}+n_{K}-y_{K})}{Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})}$
+under $H_{i}\in\{H_{C},H_{+1},\dots,H_{+K}\}$ and with
+$Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ $=\int_{0}^{1}\frac{\theta_{i}^{a_{i}-1}(1-\theta_{i})^{b_{i}-1}}{\mathrm{B}(a_{i},b_{i})}\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}I_{\theta_{i}}(a_{j},b_{j})\,\mathrm{d}\theta_{i},$
+where $I_{x}(a,b)$ is the CDF of the beta distribution.
+These marginal likelihoods can be efficiently computed with numerical integration.
+For priors with integer hyperparameters, there even exist closed-form solutions (Howard,, 1998; Miller,, 2015).
+For a specified $\Pr(H_{0})$, we may again distribute the remaining prior probability among the other hypotheses by $\Pr(H_{+i})=\{1-\Pr(H_{0})\}\times Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ to ensure that the averaged prior is again the beta prior.
+Specifying uniform priors ($a_{j}=b_{j}=1~\forall~j\in\{C,1,\dots,K\}$) is an intuitive default that ensures that each treatment receives equal prior probability.
+Plugging the marginal likelihoods and prior probabilities into (1) produces posterior probabilities, which in turn can be used to obtain randomization probabilities.
 5 Reanalyzing the ECMO trial
-The ECMO trial (Bartlett et al.,, 1985) investigated the efficacy of ECMO (extracorporeal membrane oxygenation) treatment in the critical care of newborns.
-It was the first prominent clinical trial to use a “randomized play-the-winner” (RPW) RAR design (Wei and Durham,, 1978).
-While some earlier non-randomized studies had shown a substantial treatment effect, a randomized study was required to confirm this.
-However, this presented an ethical dilemma, as the investigators were convinced that the risk of death would be much higher in the control group than in the ECMO group.
-To mitigate this, the RPW design was chosen.
-The outcome of the trial was extreme: The first newborn was randomized to receive ECMO treatment and survived.
-The second newborn was randomized to receive the control treatment and died.
-The ten subsequently enrolled newborns were all randomized to ECMO, and all survived (Bartlett et al.,, 1985).
-The trial was then stopped for efficacy and its results published.
-However, the unusual outcome sparked intense debates about whether the trial was even a proper randomized clinical trial.
-Several subsequent trials were conducted, all re-establishing the effectiveness of ECMO, which is now a standard treatment (Bartlett,, 2024).
-In the following, we will reanalyze the ECMO data.
-Figure 7: Evolution of Bayesian RAR randomization probabilities (top plot) and posterior probability of a beneficial ECMO treatment effect (bottom plot) for data from the ECMO trial.
-The top plot in Figure 7 shows Bayesian RAR probabilities computed from the ECMO data sequence.
-The normal approximation and the exact binomial methods were used, as well as the RPW method that was originally used in the ECMO trial.
+The ECMO trial (Bartlett et al.,, 1985; Bartlett,, 2024) investigated ECMO (extracorporeal membrane oxygenation) treatment in critically ill newborns.
+While earlier non-randomized studies suggested a substantial treatment effect, a randomized study was required to confirm this.
+However, investigators were convinced the mortality risk was much higher in the control group, posing an ethical dilemma.
+To mitigate this, a “randomized play-the-winner” (RPW) RAR design (Wei and Durham,, 1978) design was chosen.
+The trial’s outcome was extreme: The first newborn was randomized to ECMO and survived.
+The second was randomized to the control and died.
+All ten subsequently enrolled newborns were randomized to ECMO and survived.
+The trial was then stopped for efficacy, yet its unusual outcome sparked debate about whether it constituted a proper randomized trial.
+Later trials confirmed the efficacy of ECMO, which is now a standard treatment.
+Figure 5: Evolution of Bayesian RAR randomization probabilities (a) and posterior probability of a beneficial ECMO treatment effect (b) for data from the ECMO trial.
+Plot (c) shows the probability of observing the ECMO allocation sequence as a function of $\Pr(H_{0})$.
+Figure 5a shows RAR probabilities computed from the ECMO data.
+The normal approximation and exact binomial methods were used, as well as the RPW method originally used in the ECMO trial.
 A standard normal prior was assigned to the log odds ratio for the normal method, while uniform priors were assigned to the probabilities for the binomial method.
-For the normal method, log odds ratio estimates were computed with a Yates’ correction (i.e., adding a half to each cell) to avoid issues with zero cells.
-We can see that this correction induces a slight anomaly at the start of the study as the probabilities from the normal method slightly decrease after the first observation despite the first patient surviving the ECMO treatment.
-This does not happen for the exact method whose randomization probability increase after every additional patient, which only happens for the normal approximation after the second patient.
+Log odds ratio estimates were computed by adding a half to each cell to avoid zero cell issues.
+This correction induces a slight anomaly, as the probabilities from the normal method slightly decrease after the first newborn despite that it surived.
+This does not happen for the exact method whose probabilities increase after every additional patient.
 Furthermore, the randomization probabilities differ notably between the normal and exact methods, even at the end of the study.
 This presumably happens because the influence of the differing priors remains relatively high after observing only 12 patients.
-Comparing the randomization probabilities for different prior probabilities of the null hypothesis for each method, we can see that $\Pr(H_{0})=0$ (Thompson sampling) shows the most extreme randomization probabilities that rapidly increase to 100%, whereas $\Pr(H_{0})=1$ (equal randomization) leaves the probabilities completely static at 50%.
-In between, the randomization probabilities gradually shift from 50% to the Thompson sampling probabilities.
-The RPW probabilities (black squares) are relatively close to the probabilities from the exact method with $\Pr(H_{0})=0.5$ at the beginning of the study and become closer to the normal approximate method with $\Pr(H_{0})=0.75$ at later time points.
-The bottom plot in Figure 7 shows the corresponding posterior probability of the ECMO treatment being effective.
-Depending on the prior probability $\Pr(H_{0})$, we can see that the posterior probability at the end of the trial may be very high (e.g., $\Pr(H_{+}\mid y)=0.99$ for $\Pr(H_{0})=0$), or only moderately favoring ECMO (e.g., $\Pr(H_{+}\mid y)=0.86$ for $\Pr(H_{0})=0.5$).
-From a Bayesian perspective, stopping the trial seems thus only a sensible decision if the prior probability of equal effects was low.
-Conversely, more evidence would have been needed if the prior probability had been higher, for example, if the prior probability were $\Pr(H_{0})=0.5$, representing a priori equipoise.
+Comparing the randomization probabilities for different prior probabilities of the null hypothesis, $\Pr(H_{0})=0$ (Thompson sampling) shows the most extreme randomization probabilities that rapidly increase to 100%, whereas $\Pr(H_{0})=1$ (equal randomization) leaves the probabilities completely static at 50%.
+In between, the probabilities gradually shift from 50% to the Thompson sampling probabilities.
+The RPW probabilities (black squares) are relatively close to the exact method with $\Pr(H_{0})=0.5$ at the beginning of the study and become closer to the normal method with $\Pr(H_{0})=0.75$ at later times.
+Figure 5b shows the corresponding posterior probability of the ECMO treatment being effective.
+Depending on $\Pr(H_{0})$, the posterior probability at the end of the trial may be very high (e.g., $\Pr(H_{+}\,|\,y)=0.99$ for $\Pr(H_{0})=0$), or only moderately favoring ECMO (e.g., $\Pr(H_{+}\,|\,y)=0.86$ for $\Pr(H_{0})=0.5$).
+Stopping the trial seems thus only a sensible decision if the prior probability of no ECMO effect was low.
+Figure 5c shows the probability of observing the ECMO allocation sequence under Bayesian RAR with different $\Pr(H_{0})$.
+The probability is highest for $\Pr(H_{0})=0$ (Thompson sampling) and decreases with increasing $\Pr(H_{0})$ until $(1/2)^{12}=0.024\%$ for $\Pr(H_{0})=1$ (equal randomization).
+Interestingly, the probability under RPW is also rather low (2.8%).
+Hence, observing the ECMO allocation sequence under Bayesian RAR would only be less likely than under RPW for $\Pr(H_{0})>0.6$ (exact method) and $\Pr(H_{0})>0.55$ (normal method), respectively.
 6 Simulation study
-We conducted a simulation study to evaluate the performance of point null Bayesian RAR for different values of the prior probability of the null hypothesis $\Pr(H_{0})$, and compare it to Thompson sampling (potentially modified with burn-in periods, probability capping, power transformations) and equal randomization.
-Considered patient performance measures were the rate of successes, the rate of extreme randomization probabilities, and sample size imbalance.
-Additionally, bias and coverage were considered to evaluate the performance of rate difference point estimates and confidence intervals under RAR, while the type I error rate and power of the corresponding tests were used to quantify hypothesis testing performance under RAR.
-A binomial data-generating mechanism was used which was partially based on the simulation studies from Robertson et al., (2023), Thall and Wathen, (2007), and Wathen and Thall, (2017).
-Detailed description of the design and results of the simulation study following the structured ADEMP approach (Morris et al.,, 2019; Siepe et al.,, 2024) are provided in Appendix B.
-A supplemental website provides an interactively explorable results dashboard (https://samch93.github.io/brar/).
-Across all simulations, we observed a trade-off between patient benefit and parameter estimation / hypothesis testing operating characteristics.
-Some methods performed better in terms of patient benefit but had worse bias, coverage, type I error rate, and power while others performed better in terms of bias, coverage, type I error rate, and power but had worse patient benefit.
-This trade-off is well described in the RAR literature (Hu and Rosenberger,, 2003; Zhang and Rosenberger,, 2005; Williamson and Villar,, 2019; Robertson et al.,, 2023).
-The main result of the simulation study regarding the newly proposed method was that, under most conditions, point null Bayesian RAR with a high prior probability of the null hypothesis $\Pr(H_{0})=0.75$ showed similar operating characteristics to Thompson sampling with capped randomization probabilities at 10% and 90% and a power transformation with $c=i/(2n)$, where $i$ is the current sample size and $n$ is the maximum sample size.
-Both point null Bayesian RAR with a prior probability of $\Pr(H_{0})=0.75$ and modified Thompson sampling could mitigate some issues with ordinary Thompson sampling.
-For instance, they exhibited less negative sample size imbalance, biased parameter estimates, undercoverage, and inflated type I error rates.
-However, this improvement came at the cost of worse patient benefit performance compared to unmodified Thompson sampling.
-For instance, the mean success rate was lower, though still considerably better than equal randomization in most cases.
-Most importantly, the variability of randomization probababilitis was much reduced with $\Pr(H_{0})=0.75$, producing rarely negative imbalances (i.e., a large proportion of patients randomized to an inferior treatment).
-Setting a lower but positive prior probability than $\Pr(H_{0})=0.75$ produced operating characteristics comparable to those from less extreme modifications of Thompson sampling, such as a power transformation with $c=1/2$.
-In sum, the simulation study demonstrated that point null Bayesian RAR has comparable statistical properties to Thompson sampling with common ad hoc modifications.
+We conducted a simulation study to evaluate the performance of null hypothesis Bayesian RAR for different values $\Pr(H_{0})$, and compare it to Thompson sampling (potentially modified with burn-in periods, probability capping, power transformations), equal randomization, the Gittins index (Gittins et al.,, 2011), and Bayes upper confidence bound (UCB) (Kaufmann et al.,, 2012).
+Considered patient benefit measures were the mean rate of successes, the rate of extreme randomization probabilities, and sample size imbalance in favor of the inferior treatment.
+Bias and coverage were used to evaluate performance of response rate difference point estimates and confidence intervals under RAR, while the type I error rate and power of the corresponding tests were used to quantify hypothesis testing performance.
+A binomial data-generating mechanism was used.
+The response probability in the control group was fixed to $\theta_{C}=0.25$ and varied in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$ to assess the behavior of the methods for different effect sizes $\text{RD}_{1}=\theta_{1}-\theta_{C}$.
+In conditions with two or three treatment groups, the corresponding probabilities were set to $\theta_{2}=\theta_{3}=0.3$.
+More detailed descriptions of the design and results of the simulation study are provided in the supplement and an online results dashboard (https://samch93.github.io/brar/).
+A trade-off between patient benefit and inferential operating characteristics was observed across most settings.
+Methods that improved patient benefit through more aggressive allocation (Gittins index, Bayes UCB, and Thompson sampling) often had worse bias, coverage, type I error rate, and power, whereas more conservative methods tended to preserve inferential validity at the cost of reduced patient benefit.
+Figure 6 illustrates this trade-off for the mean success rate and coverage.
+This trade-off is well documented (Hu and Rosenberger,, 2003; Zhang and Rosenberger,, 2005; Williamson and Villar,, 2019; Robertson et al.,, 2023).
+However, there were also settings with a small “window” in which patient benefit could vary while inferential performance remained essentially unchanged.
+For example, with a burn-in of 50, two treatment groups ($K=2$), and a large effect size ($\text{RD}_{1}=0.2$), coverage was roughly constant across $\Pr(H_{0})$, whereas the mean success rate changed noticeably.
+Increasing $\Pr(H_{0})$ reduced imbalance toward the inferior treatment, though negative imbalance was also strongly influenced by effect size.
+In conditions with large effect size ($\text{RD}_{1}=0.2$), the proportion of negative imbalance was often lower for RAR methods (including Thompson sampling) than equal randomization, thus confirming the results from Robertson et al., (2023).
+These findings highlight the importance of evaluating operating characteristics during trial design, since appropriately specified RAR designs may enable improved patient benefit without substantial loss of inferential performance.
+Figure 6: Plot (a) shows the mean rate of successes (i.e., the number of successes in a study divided by its sample size averaged over all 10’000 simulation repetitions).
+The maximum Monte Carlo standard error (MCSE) is 0.051%.
+Plot (b) shows the empirical coverage rate of the 95% Wald confidence interval for the response rate difference $\text{RD}_{1}$.
+The maximum MCSE is 0.46%.
+Under most conditions, Bayesian RAR with $\Pr(H_{0})=0.75$ showed similar operating characteristics to Thompson sampling with capped randomization probabilities at 10%/90% and power transformation $c=i/(2n)$, where $i$ is the current sample size and $n$ is the maximum sample size.
+Both mitigated issues with ordinary Thompson sampling, exhibiting less negative sample size imbalance, less bias, improved coverage, and reduced type I error rates.
+This came at the cost of worse patient benefit performance, for instance, lower mean success rate, though still better than equal randomization in most cases.
+Smaller values of $\Pr(H_{0})$ produced operating characteristics comparable to less extreme modifications, such as $c=1/2$.
+Finally, the results were similar for exact binomial and approximate normal versions of the method (supplement), suggesting some robustness to prior and model choice.
+While these results demonstrated that null hypothesis Bayesian RAR has comparable properties to capped Thompson sampling for realistic trial sizes, it is important to note that the two are expected to differ asymptotically.
+Capped Thompson sampling enforces bounds on randomization probabilities, whereas uncapped Bayesian RAR probabilities can converge to 100% or 0%.
+The observed similarity is thus only a finite-sample phenomenon.
 7 Discussion
-In this paper we have proposed a modification of standard Bayesian RAR (Thompson sampling) via recasting of the problem in a Bayesian hypothesis testing framework and the introduction of a point null hypothesis.
-While the plausibility of point null hypotheses is often a matter of philosophical debate (see e.g., Berger and Delampady,, 1987; Ly and Wagenmakers,, 2022), this method is useful in the RAR setting, as it can interpolate between equal randomization and Thompson sampling by changing the prior probability of the null hypothesis $\Pr(H_{0})$.
-This allows experimenters to balance patient benefit with classical operating characteristics, such as power, type I error rate, bias, and coverage.
-For large values of $\Pr(H_{0})$, we observed behaviors and operating characteristics similar to those obtained with ad hoc modifications of Thompson sampling, such as capping, burn-ins, and power transformations.
-Our method is implemented in the free and open source R package brar for binomial outcomes and for data summarized by approximately normal effect estimates.
-The latter makes the method applicable to many settings, for example, settings where treatment effects are estimated with regression analyses.
-One advantage of our framework is that the randomization probabilities coherently correspond with the available statistical evidence (in the form of Bayes factors) and beliefs (in the form of posterior probabilities).
-In principle, both could also be used as decision-making tools instead of relying on frequentist test criteria.
-For instance, if the posterior probability of a treatment’s superiority is greater than 0.99, say, a study could be stopped.
-This also makes sense from the perspective that it is unnatural to randomize participants with extreme randomization probabilities that are associated with such high posterior probabilities.
-Although we conducted a simulation study to understand the method’s basic behavior, more realistic and comprehensive evaluations are needed to understand its applicability in real-world conditions.
-For example, the method needs to be evaluated in combination with futility stopping (e.g., dropping of treatment arms which are shown to be ineffective at an interim analysis).
-Another issue to consider is how to select the prior probability of the null hypothesis.
-In our simulation study, we found that setting a value of $\Pr(H_{0})=0.75$ mitigated many of the issues with Thompson sampling.
-However, other choices could be considered, such as setting a higher value.
-Similar considerations apply to the prior distributions of the parameters, as we did not assess the effect of varying these on the operating characteristics.
-For instance, rather than setting the correlation of the multivariate normal prior to achieve equal randomization probabilities, it could be beneficial to set it so that the prior probability of the control being superior is always $\Pr(H_{-})=0.5$ and the remaining probability is distributed equally among the treatments.
-Future work may therefore investigate whether a more efficient RAR procedure can be obtained by specifying a certain prior distribution.
-Finally, in many clinical trial settings, RAR methods are not directly applicable because outcomes such as death may only be observed after long follow-up periods, by which time recruitment and randomized allocation will already have finished.
-An alternative could be to perform point null Bayesian RAR with an informative surrogate outcome that is sooner observed than the primary outcome (Gao et al.,, 2024).
-Acknowledgments
-We thank František Bartoš and Małgorzata Roos for valuable comments on drafts of the manuscript. The acknowledgment of these individuals does not imply their endorsement of the paper.
-Conflict of interest
-We declare no conflict of interest.
-Software and data
-Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/brar.
-A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.17248628.
-We used the statistical programming language R version 4.5.0 (2025-04-11) for analyses (R Core Team,, 2024) along with the ggplot2 (Wickham,, 2016), dplyr (Wickham et al.,, 2023), SimDesign (Chalmers and Adkins,, 2020), mvtnorm (Genz and Bretz,, 2009), ggh4x (van den Brand,, 2024), ggpubr (Kassambara,, 2023), and knitr (Xie,, 2015) packages.
+We have proposed a modification of Thompson sampling by recasting the problem in a Bayesian hypothesis testing framework and introducing a null hypothesis of equal effectiveness.
+This method can interpolate between equal randomization and Thompson sampling by changing the prior probability of the null hypothesis $\Pr(H_{0})$.
+This allows balancing patient benefit with inferential properties.
+For large values of $\Pr(H_{0})$, we observed similar characteristics as with ad hoc modifications of Thompson sampling.
+One advantage of the method is that randomization probabilities coherently correspond to the available statistical evidence (in the form of Bayes factors) and beliefs (in the form of posterior probabilities).
+Both could also serve as decision-making tools instead of frequentist criteria.
+For example, a study could be stopped if the posterior probability of efficacy exceeds 0.99.
+This also aligns with the perspective that it is unnatural to randomize patients using extreme randomization probabilities associated with such high posterior probabilities.
+Although we conducted a simulation study to understand the method’s basic behavior, more evaluations are needed to understand its real-world applicability.
+For example, the method needs to be evaluated in combination with futility stopping (e.g., dropping of arms at interim) and under time trends (e.g., effects attenuating over time because the standard of care improves).
+RAR is often impractical when primary outcomes require long follow-up.
+An alternative could be to perform Bayesian RAR with a surrogate outcome that is sooner observed (Gao et al.,, 2024).
+Finally, further theoretical work is needed to understand the asymptotic behavior of the method.
+Most importantly, the empirically observed convergence of randomization probabilities to 50% under $H_{0}$ remains to be shown rigorously.
+Regarding prior choice, $\Pr(H_{0})=0.75$ mitigated many Thompson sampling issues in our study, yet higher values may also be considered.
+Similar considerations apply to the prior distributions of the parameters.
+For instance, rather than setting the correlation of the multivariate normal prior to achieve equal randomization probabilities, it could be set so that the prior probability of the control being superior is always $\Pr(H_{0})=0.5$ and the remaining probability is equally distributed among the treatments.
+Similarly, concurrent allocation to the control group may be desirable (Trippa et al.,, 2012).
+Future work may therefore investigate whether more efficient RAR can be obtained by tuning the prior distribution.
+Acknowledgments and conflict of interest
+We thank František Bartoš and Małgorzata Roos for valuable comments. We thank the associate editor and two anonymous reviewers for excellent suggestions. We declare no conflict of interest.
+Supplementary materials
+Supplementary materials and data and code are available with this paper at the Biometrics website on Oxford Academic.
+Data availability
+Code and data to reproduce our results are available at https://github.com/SamCH93/brar and archived at https://doi.org/10.5281/zenodo.17248628.
 Appendix A The R package brar
-Our R package can be installed by running remotes::install_github(repo = "SamCH93/brar", subdir = "package") in an R session (requires the remotes package, which is available on CRAN).
-The main functions of the package are brar_normal and brar_binomial, which implement the approximate normal method from Section 3 and the exact binomial method from Section 4.
+Our R package can be installed by running install.packages("brar") in an R console.
+The main functions of the package are brar_normal and brar_binomial, which implement the normal method from Section 3 and the binomial method from Section 4 in the main text.
 The following code chunk illustrates how the latter function can be used.
@@
 […]
-Appendix B Simulation study
+Appendix B Null hypothesis Bayesian RAR under multivariate normality
+Suppose there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to the control.
+A natural generalization univariate normal Bayesian RAR is to stack the estimates into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume a $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\mid\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the (assumed to be known) covariance matrix of $\boldsymbol{\hat{\theta}}$.
+For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ its estimated covariance matrix.
+The hypotheses from Section 2 in the main text then translate into
+$H_{-}\colon\max\{\theta_{1},\dots,\theta_{K}\}<0$ vs. $H_{0}\colon\theta_{1}=\dots=\theta_{K}=0$ vs. $H_{+i}\colon\theta_{i}=\max\{\theta_{1},\dots,\theta_{K}\}>0$
+for $i=1,\dots,K$.
+All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution.
+In analogy to the one treatment case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of the corresponding hypothesis (e.g., for hypothesis $H_{+i}$, the space in $\mathbb{R}^{K}$ where the $i$th component is positive and larger than and all other components).
+As in the one group case, it seems a sensible default to center the prior on $\boldsymbol{0}$ to encode clinical equipoise.
+Similarly, the prior hypothesis probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ may again be specified so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses.
+Additionally, specifying a prior covariance matrix $\boldsymbol{\mathcal{T}}$ with uniform correlation of 0.5 ensures equal treatment prior probabilities for any number of treatment groups $K$, and thus seems a sensible default in the absence of prior knowledege.
+The normal-normal conjugate framework allows us to derive marginal likelihoods in closed-form. The marginal likelihood under $H_{0}$ is
+$p(\boldsymbol{\hat{\theta}}\mid H_{0})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{0},\boldsymbol{\Sigma}),$
+while the marginal likelihood under $H_{-}$ is
+$p(\boldsymbol{\hat{\theta}}\mid H_{-})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu}_{*},\boldsymbol{\mathcal{T}}_{*})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{\mu},\boldsymbol{\mathcal{T}})}$
+with $\mathrm{N}_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ and $\Phi_{K}(\boldsymbol{x}\mid\boldsymbol{m},\boldsymbol{V})$ the density and cumulative distribution functions of the $K$-variate normal distribution with mean vector $\boldsymbol{m}$ and covariance matrix $\boldsymbol{V}$ evaluated at $\boldsymbol{x}$, and posterior mean $\boldsymbol{\mu}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}(\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\theta}}+\boldsymbol{\mathcal{T}}^{-1}\boldsymbol{\mu})$ and covariance $\boldsymbol{\mathcal{T}}_{*}=(\boldsymbol{\Sigma}^{-1}+\boldsymbol{\mathcal{T}}^{-1})^{-1}$.
+Finally, the marginal likelihood under $H_{+i}$ is given by
+$p(\boldsymbol{\hat{\theta}}\mid H_{+i})$ $=\mathrm{N}_{K}(\boldsymbol{\hat{\theta}}\mid\boldsymbol{\mu},\boldsymbol{\Sigma}+\boldsymbol{\mathcal{T}})\times\frac{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu}_{*},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}_{*}\boldsymbol{A}_{i,K}^{\top})}{\Phi_{K}(\boldsymbol{0}\mid\boldsymbol{A}_{i,K}\boldsymbol{\mu},\boldsymbol{A}_{i,K}\boldsymbol{\mathcal{T}}\boldsymbol{A}_{i,K}^{\top})}$
+where $\boldsymbol{A}_{i,K}$ is a $K\times K$ contrast matrix that maps $\boldsymbol{\theta}$ to the space where the negative orthant corresponds to the space of hypothesis $H_{+i}$.
+For example, for $i=2$ and $K=3$, the matrix is
+$\boldsymbol{A}_{2,3}=\begin{bmatrix}0&-1&0\\ 1&-1&0\\ 0&-1&1\\ \end{bmatrix}$
+with the first row encoding the constraint of $\theta_{2}$ being positive, and the second and third rows encoding the constraints of $\theta_{2}$ being larger than $\theta_{1}$ and $\theta_{3}$, respectively.
+As expected, for $K=1$, these marginal likelihoods reduce to the univariate ones (7) in the main text.
+Moreover, since they are all available in closed-form, Bayes factors, posterior probabilities, and randomization probabilities are also available in closed-form.
+Crucially, only the exact evaluation of multivariate normal densities and cumulative distribution functions is required, and both are efficiently implemented in statistical software (e.g., in the mvtnorm R package, Genz and Bretz,, 2009).
+This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve additional complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
+Figure S1: Evolution of Bayesian RAR probabilities for 3 treatments and a control group.
+In each step, one observation from each group is simulated from a normal distribution with a standard deviation of 1 and corresponding mean differences as indicated in the panel titles (treatment 1 is the most effective treatment), and a mean difference effect estimate $\boldsymbol{\hat{\theta}}$ and its estimated covariance matrix $\boldsymbol{\Sigma}$ computed with linear regression.
+Randomization probabilities are then computed using a normal spike-and-slab prior for centered at the origin and with covariance matrix $\boldsymbol{\mathcal{T}}$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, and for different prior probabilities $\Pr(H_{0})$.
+Figure S1 shows sequences of randomization probabilities computed from simulated normal data with $K=3$ treatment groups.
+We see that assigning a prior probability $\Pr(H_{0})=1$ leads to static equal randomization at $\pi_{i}=1/(K+1)=25\%$, while assigning $\Pr(H_{0})=0$ (Thompson sampling) leads to the most variable randomization probabilities.
+Since data are simulated assuming that treatment 1 is the most effective, randomization probabilities based on $\Pr(H_{0})<1$ converge towards $\pi_{1}=100\%$ and toward $0\%$ for the remaining treatments.
+While convergence is the fastest for $\Pr(H_{0})=0$, this prior probability also accidentally produces rather high randomization probabilities for the control group at the start of the study, which is less pronounced for positive prior probabilities $\Pr(H_{0})>0$.
+Appendix C Simulation study
 We now describe the design and results of our simulation study following the structured ADEMP approach (Morris et al.,, 2019; Siepe et al.,, 2024).
@@
 A website with additional details and results is provided at https://samch93.github.io/brar/.
-B.1 Aims
-The aim of the simulation study is to evaluate the design characteristics of the newly proposed point null Bayesian RAR approach, and compare it to existing methods.
-B.2 Data-generating mechanism
+C.1 Aims
+The aim of the simulation study is to evaluate the design characteristics of the newly proposed null hypothesis Bayesian RAR approach, and compare it to existing methods.
+C.2 Data-generating mechanism
 The data-generating mechanism was inspired by the simulation studies from Robertson et al., (2023), Thall and Wathen, (2007), and Wathen and Thall, (2017).
@@
 Parameters were chosen similar to the simulation study from Robertson et al., (2023).
-We vary the sample size $n\in\{200,654\}$ to represent low and high powered studies, the number of treatment groups $K\in\{1,2,3\}$, and probability in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$.
+We vary the sample size $n\in\{200,654\}$ to represent low and high powered studies, the number of treatment groups $K\in\{1,2,3\}$, and the probability in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$.
 The probability in the control group and the remaining groups is always fixed at $\theta_{C}=0.25$ and $\theta_{2}=\theta_{3}=0.3$, respectively.
@@
 These come with additional parameters that are, however, considered as method tuning parameters rather than true underlying parameters.
-B.3 Estimands and other targets
+C.3 Estimands and other targets
 The primary interest of this simulation study lies in assessing the patient benefit characteristics of different RAR methods.
-Additionally, the estimand of interest is the rate difference $\text{RD}_{1}=\theta_{1}-\theta_{C}$ and the target of interest is the null hypothesis of $\text{RD}_{1}=0$.
-B.4 Methods
-We consider the above described point null RAR methods.
+Additionally, the estimand of interest is the response rate difference $\text{RD}_{1}=\theta_{1}-\theta_{C}$ and the target of interest is the null hypothesis of $\text{RD}_{1}=0$.
+C.4 Methods
+We consider the null hypothesis Bayesian RAR method described in the main text.
 The prior probability of $H_{0}$ is a tuning parameter and controls the variability of the randomization probabilities.
@@
 This re-normalization is only performed for randomization probabilities greater than 10%, as these would otherwise be reduced again to probabilities less than 10%.
-In case, a re-normalized probability becomes less than 10%, it is also capped at 10% and second re-normalization performed.
-Finally, for equal randomization ($\Pr(H_{0})=1$), no burn-in, capping, or power transformation conditions are simulated as these manipulations have no effect.
-B.5 Performance measures
-Different performance measures were used. Patient benefit was quantified with:
+In case, a re-normalized probability becomes less than 10%, it is also capped at 10% and a second re-normalization performed.
+For equal randomization ($\Pr(H_{0})=1$), no burn-in, capping, or power transformation conditions are simulated as these manipulations have no effect.
+Finally, upon suggestions from the associate editor, we also include two methods from the multi-armed bandit literature in the simulation study to compare null hypothesis Bayesian RAR to more aggressive allocation methods – the Gittins index (Gittins et al.,, 2011; Villar et al.,, 2015) and the Bayesian upper confidence bound (Bayes UCB) methods (Kaufmann et al.,, 2012).
+For the Gittins index, we use the implementation from the RARtrials R package (Xu et al.,, 2025) with a discount factor of $d=0.995$.
+For Bayes UCB we use uniform priors for each probability along with the tuning parameter $c=0$, as recommended in Kaufmann et al., (2012) based on their simulation results.
+C.5 Performance measures
+Patient benefit was quantified with:
   • The mean rate of successes per study
@@
 For $K=1$, this reduces to the $\hat{S}_{0.1}$ sample size imbalance measure from Robertson et al., (2023), which in turn was inspired by the performance evaluation in the simulation study from Thall et al., (2015).
+Note that the $\hat{S}_{0.1}$ measure is difficult to translate into an actual patient benefit measure because the patient benefit loss related to imbalance depends on the size of the treatment effect (Robertson et al.,, 2023).
+For instance, imbalance is less consequential if the treatment effect is small than if it is large.
+Nevertheless, we include the $\hat{S}_{0.1}$ measure for comparability with previous simulation studies.
 Parameter estimation and hypothesis testing performance was quantified with:
@@
 MCSEs were calculated using the formulae from Siepe et al., (2024) and are provided for all measures in the following figures and the supplemental website.
-B.6 Computational aspects
-The simulation study was run on a server running Debian GNU/Linux 13 (trixie) and R version 4.5.0 (2025-04-11).
+C.6 Computational aspects
+The simulation study was run on a server running Debian GNU/Linux forky/sid and R version 4.5.2 (2025-10-31).
 The SimDesign R package was used to organize and run the simulation study (Chalmers and Adkins,, 2020).
 Our newly developed brar R package was used to perform Bayesian RAR.
-Code and data to reproduce this simulation study are available at https://github.com/SamCH93/brar.
-B.7 Results
+Code and data to reproduce this simulation study are available at https://github.com/SamCH93/brar and persistenly archived at https://doi.org/10.5281/zenodo.17248628.
+C.7 Results
 Convergence
@@
 The mean rate of successes is shown in Figure LABEL:fig:successrate.
-It was generally the highest for Thompson sampling and lowest for equal randomization.
+It was generally the highest for the Gittins index, followed by Bayes UCB, Thompson sampling, and being the lowest for equal randomization.
 The normal approximation and exact method produced mostly similar rates, with the normal method sometimes showing slightly higher rates (e.g., for $K=3$ and $\text{RD}_{1}=0.2$).
@@
 This is also visible in Figure LABEL:fig:extreme where more extreme randomization probabilities are observed with increasing sample size and rate difference for all methods but the ones with capping.
+Finally, Bayes UCB and the Gittins index methods always produced extreme randomization probabilities, as these methods deterministically allocate the next patient to the group with the highest upper confidence bound for and Gittins index, respectively.
 Negative sample size imbalance
 Figure LABEL:fig:imbalance shows negative sample size imbalance as quantified by the $\hat{S}_{0.1}$ metric.
-Negative imbalance was the greatest for Thompson sampling and reduced when modifications were applied.
+The Gittins index generally showed the highest negative imbalance, followed by Thompson sampling and Bayes UCB.
+Among the methods with $\Pr(H_{0})<1$, negative imbalance was the greatest for Thompson sampling and reduced when modifications were applied.
 Similarly, increasing the prior probability of $H_{0}$ decreased negative imbalance, in some cases even below Thompson sampling with modifications.
+As in the simulation study from Robertson et al., (2023), negative imbalance of Thompson sampling and Bayesian RAR with $\Pr(H_{0})<1$ was higher than under equal randomization in conditions with small effect size $\text{RD}_{1}=0.1$, but equal or lower than under equal randomization in conditions with large effect size $\text{RD}_{1}=0.2$.
 Bias and coverage
@@
 In conditions where there was no difference ($\text{RD}_{1}=0$), all methods produced unbiased point estimates, though all methods but equal randomization also showed undercoverage.
-Bias occurred in conditions with non-zero rate differences, the bias being the greatest for Thompson sampling and decreasing to some extent when modifications were introduced or the prior probability of $H_{0}$ increased.
-Similarly, modifications or increasing prior probabilitities improved coverage, although they still remained suboptimal in most conditions.
-For large rate difference conditions, coverage was much better for the normal than the exact version of RAR.
-Similarly, bias was in some cases larger for the exact compared to the normal version.
+Bias occurred in conditions with non-zero rate differences, the bias being the greatest for the Gittins index, Bayes UCB, and Thompson sampling, and decreasing to some extent when modifications were introduced or the prior probability of $H_{0}$ increased.
+Similarly, modifications or increasing prior probabilitities improved coverage, although coverage still remained suboptimal in most conditions.
+For large rate difference conditions, coverage was often better for the normal than the exact version of RAR.
+Similarly, bias was in some cases somewhat larger for the exact compared to the normal version.
+This could potentially be due to the normal prior distribution on the log odds ratio being better tuned to effect sizes in the data-generating mechanism difference compared to the uniform priors on the probabilities under the binomial model.
 Type I error rate and power
 Figures LABEL:fig:t1e and LABEL:fig:power show empirical type I error rate and power associated with the Wald test of $\text{RD}_{1}=0$.
-We see that standard Thompson sampling shows an inflated type I error rate above the nominal 2.5% which is reduced to some extent by increasing either the prior probability of $H_{0}$ or applying modifications.
-In the same way, Thompson sampling exhibits reduced power compared to equal randomization, which is again alleviated by modifications or positive prior probability of the null hypotheses.
+We see that the Gittins index, Bayes UCB, and Thompson sampling methods showed an inflated type I error rate above the nominal 2.5% which was reduced to some extent by increasing either the prior probability of $H_{0}$ or applying modifications.
+In the same way, Gittins index, Bayes UCB, and Thompson sampling often exhibited reduced power compared to equal randomization, which was again alleviated by modifications or increasing $\Pr(H_{0})$.
 In small sample sizes and large rate differences, power was slightly increased for the exact compared to the normal version of RAR, while for Thompson sampling the type I rate was slightly higher for the exact compared to the normal version.
+Appendix D Software and data
+Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/brar.
+A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.17248628.
+We used the statistical programming language R version 4.5.2 (2025-10-31) for analyses (R Core Team,, 2024) along with the ggplot2 (Wickham,, 2016), dplyr (Wickham et al.,, 2023), SimDesign (Chalmers and Adkins,, 2020), mvtnorm (Genz and Bretz,, 2009), ggh4x (van den Brand,, 2024), ggpubr (Kassambara,, 2023), RARtrials (Xu et al.,, 2025), and knitr (Xie,, 2015) packages.